### v7

Abstract
We study the bias of classical quantile regression and instrumental variable quantile regression estimators. While being asymptotically first-order unbiased, these estimators can have non-negligible second-order biases. We derive a higher-order stochastic expansion of these estimators using empirical process theory. Based on this expansion, we derive an explicit formula for the second-order bias and propose a feasible bias correction procedure that uses finite-difference estimators of the bias components. The proposed bias correction method performs well in simulations. We provide an empirical illustration using Engel’s classical data on household food expenditure.
JEL Classification: C21, C26.
Keywords: instrumental variables, higher-order stochastic expansion, Bahadur-Kiefer expansion, finite-difference estimators, mixed integer linear programming (MILP), Engel curve.
1 Introduction
Many interesting empirical applications of classical quantile regression (QR) (Koenker and Bassett, 1978) and instrumental variable quantile regression (IVQR) (Chernozhukov and Hansen, 2005, 2006) feature small sample sizes, which can arise either as a result of a limited number of observations or when estimating tail quantiles, or both (e.g., Chernozhukov, 2005; Elsner et al., 2008; Chernozhukov and Fernández-Val, 2011; Adrian and Brunnermeier, 2016; Adrian et al., 2019). QR and IVQR estimators are nonlinear and can thus exhibit substantial biases in small samples. In this paper, we theoretically characterize these biases and develop a feasible bias correction procedure.
To study the biases, we start by deriving a higher-order stochastic expansion of the classical QR and exact IVQR estimators. [^2] We define exact IVQR estimators as estimators that exactly minimize a norm of the sample moment conditions. Such estimators can be obtained using mixed-integer programming (MIP) methods (e.g., Chen and Lee, 2018; Zhu, 2019). See Appendix D. Such an expansion is needed because the higher-order terms contribute nonzero biases while the first-order term does not. We derive explicit expressions and uniform (in the quantile level) rates for the components in the expansion, building on the empirical process arguments of Ota et al. (2019). This expansion can be thought of as a refined Bahadur-Kiefer (BK) representation of the estimator that decomposes the nonlinear component into terms up to the order $O_{p}\left(n^{-1}\right)$ and a $O_{p}\left(n^{-5/4}\sqrt{\log n}\right)$ remainder. We also derive a uniform BK representation for generic IVQR estimators after a feasible 1-step Newton correction (see Appendix E).
Using the stochastic expansion, we study the bias of QR and exact IVQR estimators. We derive a bias formula based on the leading terms up to order $O_{p}\left(n^{-1}\right)$ in the expansion, which we refer to as the second-order (asymptotic) bias. This approach of focusing on the moments of the leading terms in the stochastic expansions is standard in the literature. [^3] See, for example, Nagar (1959); Newey and Smith (2004); Kato et al. (2012); Galvao and Kato (2016); Kaplan and Sun (2017); Hahn et al. (2023), among others. The second-order bias formula provides an approximation of the actual bias that yields a feasible correction. Our results explicitly account for the $O\left(n^{-1}\right)$ (up to a logarithmic factor) bias due to nonzero sample moments at the estimator. Our proof strategy is different from the generalized function heuristic used in the existing literature (Phillips, 1991; Lee et al., 2017, 2018), which does not account for all the leading terms up to $O\left(n^{-1}\right)$ in the bias (see Section 3.2 and Appendix C for further discussion and examples). The missing terms can be important bias contributors in practice, as we document in the empirical application in Section 5.
A feasible bias correction procedure then follows from the second-order bias formula. We propose finite-difference estimators of all the components in the formula. These estimators admit higher-order expansions that allow us to select bandwidth rates. In particular, our finite-difference estimator for the Jacobian coincides with Powell (1986)’s classical estimator, and the bandwidth rate we derive coincides with the AMSE optimal bandwidth choice in Kato (2012). We show that the resulting (analytically) bias-corrected estimator has zero second-order bias. This result is in contrast to the commonly-used bootstrap bias correction approaches (Horowitz, 2001), which may not capture the higher-order bias terms of quantile estimators (see Knight, 2003).
We evaluate the performance of our bias correction procedure in a Monte Carlo simulation study. The simulations show that the theoretical (infeasible) bias formula describes well the second-order bias of classical QR and exact IVQR. We find that the proposed feasible bias correction can effectively reduce the bias in many cases. The gains from bias correction are particularly prominent in settings with large bias such as for the IVQR estimators under endogeneity. The impact of the bias correction on the root MSE (RMSE) is rather small and ambiguous.
We illustrate the bias correction approach by revisiting the relationship between food expenditure and income based on the original Engel (1857) data (e.g., Koenker and Bassett, 1982; Koenker and Hallock, 2001). [^4] The QR approach was used more recently to analyze changes in heterogeneity in the income elasticity over time of various consumer spending categories in Taylor and Houthakker (2009). Our results highlight the importance of bias correction in empirical applications with small sample sizes. Specifically, we find that the second-order bias of classical QR is empirically relevant: it can be larger than 50% of the standard error, which is substantial given that this is the second-order bias.
Roadmap.
The remainder of the paper is organized as follows. Section 2 describes the model and the estimators. Section 3 provides our main theoretical results. Section 4 presents the Monte Carlo simulation results. Section 5 contains the empirical application. Section 6 concludes. All the proofs and some additional details are given in the Appendix.
2 Model and estimators
Consider a setting with a continuous outcome variable $Y$, a $(k\times 1)$ vector of covariates $W$, and a $(k\times 1)$ vector of instruments $Z$. We assume throughout that $k$ is fixed. Every observation $(Y_{i},W_{i},Z_{i})$, $i=1,\dots,n$, is jointly drawn from a distribution $P$. We assume that $(Y_{i},W_{i},Z_{i})$ is i.i.d., and we will sometimes suppress the index $i$ to lighten up the notation. The parameter of interest $\theta_{\tau}\in\Theta\subset\mathbb{R}^{k}$ is defined as a solution to the following unconditional quantile moment restrictions,
$\mathbb{E}[(1\{Y\leq W^{\prime}\theta_{\tau}\}-\tau)Z]=0,\quad\tau\in(0,1).$ (1)
We consider two cases: (i) classical QR, where $Z=W$ (Koenker and Bassett, 1978), and (ii) linear IVQR, where $Z\neq W$ in general (Chernozhukov and Hansen, 2006, 2008).
The classical QR estimator of $\theta_{\tau}$ is a solution to the following convex minimization problem,
$\hat{\theta}_{\tau,QR}\in\operatorname{argmin}_{\theta\in\Theta}\mathbb{E}_{n}\rho_{\tau}(Y-W^{\prime}\theta),$ (2)
where $\rho_{\tau}(u)=u(\tau-1\{u<0\})$ is the check function (Koenker, 2005) and $\mathbb{E}_{n}$ denotes the sample average, i.e., the expectation with respect to the empirical measure. For IVQR, we consider estimators that exactly minimize the $p$-norm of the sample moments,
$\hat{\theta}_{\tau,p}\in\operatorname{argmin}_{\theta\in\Theta}||\hat{g}_{\tau}(\theta)||_{p},$ (3)
where $p\in[1,\infty]$ and $\hat{g}_{\tau}(\theta)\triangleq\mathbb{E}_{n}(1\{Y\leq W^{\prime}\theta\}-\tau)Z$. This class of exact IVQR estimators includes GMM, which corresponds to $p=2$ as in Chen and Lee (2018) for just-identified models, and the estimator proposed by Zhu (2019), which corresponds to $p=\infty$. The cases $p=1$ and $p=\infty$ have computationally convenient mixed integer linear programming (MILP) representations, while the MILP formulation for $p=2$ has many more decision variables. In our Monte Carlo simulations, we use $p=1$ for computational convenience (see Appendix D).
We use the notation $g_{\tau}(\theta)\triangleq\mathbb{E}(1\{Y\leq W^{\prime}\theta\}-\tau)Z$ for the unconditional moment restrictions as a function of $\theta\in\Theta$, and write $G(\theta)\triangleq\partial_{\theta}\mathbb{E}Z1\{Y\leq W^{\prime}\theta\}=\partial_{\theta}g_{\tau}(\theta)$ for its derivative. We maintain the following standard identification assumptions.
Assumption 1 (Identification).
  1. 1.
    $\theta_{\tau}$ is the unique solution to $g_{\tau}(\theta)=0$ over a compact set $\Theta\subset\mathbb{R}^{k}$, and $\theta_{\tau}$ is in the interior of $\Theta$ for all $\tau\in[\varepsilon,1-\varepsilon]$ for some $\varepsilon>0$.
  2. 2.
    The Jacobian $G(\theta_{\tau})$ has full rank for all $\tau\in(0,1)$.
As noted by Chernozhukov and Hansen (2006), “compactness [of the parameter space $\Theta$] is not restrictive in micro-econometric applications” (p.502). Throughout the paper, we use the short notation $G$ for $G(\theta_{\tau})$ whenever it does not lead to ambiguity.
We impose the following smoothness assumptions on the conditional density and its derivatives. Such assumptions are standard in the literature on higher-order properties of quantile estimators (e.g., Ota et al., 2019).
Assumption 2 (Conditional density).
The conditional density of $Y_{i}$ given $(W_{i},Z_{i})$, $f_{Y}(y|w,z)$, exists, is a.s. three times continuously differentiable on $supp(Y)$, and there exists a constant $\bar{f}$ such that $|f^{(r)}_{Y}(y|w,z)|\leq\bar{f}$ for all $(y,w,z)\in supp(Y)\times supp(W)\times supp(Z)$, where $r=0,1$ and $f_{Y}^{(r)}(\cdot|w,z)$ is the $r$-th derivative of $f_{Y}(\cdot|w,z)$.
In our theoretical analysis of the bias, we will often work with a related object, the conditional density $f_{\varepsilon_{\tau}}(e|W,Z)\triangleq f_{Y}(e+W^{\prime}\theta_{\tau}|W,Z)$ of the quantile residual $\varepsilon_{\tau}\triangleq Y-W^{\prime}\theta_{\tau}$.
Finally, we assume that the regressors and the instruments have bounded higher-order moments.
Assumption 3 (Regressors and instruments).
There exists constants $m<\infty$ and $\gamma\geq 6$ such that $\mathbb{E}|W_{j}|^{\gamma}\leq m$ and $\mathbb{E}|Z_{j}|^{\gamma}\leq m$ for all $j=1,\dots,k$.
Assumption 3 guarantees the existence of all relevant moments of the terms involving $W$ and $Z$ in the higher-order derivatives of the moment conditions and the bias correction (e.g., $\mathbb{E}Z_{\ell}W_{j}W_{q}W_{r}$). The power parameter $\gamma$ (as we show below) determines the upper bound on the rate at which the norm of the sample moment functions converges to zero — higher $\gamma$ implies faster convergence to zero.
3 Asymptotic theory for bias correction
To derive a bias correction procedure, we follow the approach of Nagar (1959) and focus on the bias of the leading terms in the asymptotic stochastic expansion of the estimator.
3.1 Stochastic expansion of quantile regression estimators
The classical first-order asymptotic theory for quantile regression estimators (e.g., Koenker and Bassett, 1978; Angrist et al., 2006; Chernozhukov and Hansen, 2006; Kaplan and Sun, 2017; Kaido and Wüthrich, 2021) is based on the following leading term,
$\hat{\xi}_{\tau}\triangleq\theta_{\tau}-G^{-1}(\theta_{\tau})\mathbb{E}_{n}Z\left(1\{Y\leq W^{\prime}\theta_{\tau}\}-\tau\right).$ (4)
For correctly specified models, $\hat{\xi}_{\tau}$ is an infeasible unbiased estimator of $\theta_{\tau}$. However, because feasible quantile estimators are nonlinear, they generally have a nonzero higher-order bias. The following theorem provides a characterization of the terms in a stochastic expansion of $\hat{\theta}_{\tau}$ up to order $O_{p}\left(n^{-1}\right)$ (ignoring logarithmic terms).
To state the result, we introduce some additional notation. For $\theta\in\Theta$, define the auxiliary functions $g^{\circ}(\theta)\triangleq\mathbb{E}1\{Y\leq W^{\prime}\theta\}Z$, $B^{\circ}_{n}(\theta)\triangleq\sqrt{n}(\mathbb{E}_{n}1\{Y\leq W^{\prime}\theta\}Z-g^{\circ}(\theta))$, and $B_{n}(\theta)\triangleq B^{\circ}_{n}(\theta)-B^{\circ}_{n}(\theta_{\tau})$. [^5] The processes $B_{n}(\theta)$ and $B^{\circ}_{n}(\theta)$ take values in the space $\ell^{\infty}(\Theta)$ of bounded functions on $\Theta$. Also, denote the Hessian of the $j$-th moment function $g_{j}$ by $\partial_{\theta}G_{j}(\theta)\triangleq\partial_{\theta}\partial_{\theta}g_{j}(\theta)$. Finally, for any $x\in\mathbb{R}^{k}$, denote by $x^{\prime}\partial_{\theta}G(\theta)x$ the vector with components $x^{\prime}\partial_{\theta}G_{j}(\theta)x$, $j=1,\dots,k$.
Theorem 1.
Suppose that Assumptions 13 hold. Consider $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ obtained from program (3) for some $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,QR}$. Then
$\hat{\theta}_{\tau}=\hat{\xi}_{\tau}+G^{-1}(\theta_{\tau})\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\frac{B_{n}(\hat{\theta}_{\tau})}{\sqrt{n}}-\frac{1}{2}(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})\right]+R_{n,\tau},$
where
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau})\|$ =$=\begin{cases}O_{p}\left(\frac{1}{n}\right),\text{ if }\hat{\theta}_{\tau}=\hat{\theta}_{\tau,\text{QR}},\\ O_{p}\left(\frac{\log n}{n^{1-\frac{2}{\gamma}}}\right),\text{ if }\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p},\end{cases}$ (5)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|B_{n}(\hat{\theta}_{\tau})\|$ $=O_{p}\left(\frac{\sqrt{\log n}}{n^{1/4}}\right),$ (6)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\xi}_{\tau}-\theta_{\tau}\|$ $=O_{p}\left(\frac{1}{n^{1/2}}\right),$ (7)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|R_{n,\tau}\|$ $=O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right).$
The proof of Theorem 1 builds on the empirical process arguments in Lemma 3 of Ota et al. (2019) and uses the maximal inequality in Corollary 5.1 of Chernozhukov et al. (2014).
The rates in Theorem 1 are uniform in the quantile level $\tau$. Uniformity is important in theory and practice because QR and IVQR methods are particularly powerful when used to analyze the entire quantile process.
Classical QR and linear IVQR are motivated by the linearity of the true conditional quantile function. In practice, linearity may be restrictive, especially when it is imposed at each $\tau\in(0,1)$. While the result in Theorem 1 remains valid if linearity fails, the interpretation is more complicated in this case. Specifically, if the true conditional quantiles are nonlinear, then the expansion in Theorem 1 applies to the pseudo-true value defined via moment condition (1). For classical QR, this pseudo-true value can be interpreted as the minimizer of a weighted mean-squared error loss function (Angrist et al., 2006). Note further that the result in Theorem 1 holds for every fixed $\tau\in(0,1)$. Thus, if linearity holds at a given $\tau$, we can interpret the expansion at this $\tau$ under the correct specification without requiring linearity at the other quantile levels.
The expansion in Theorem 1 can be thought of as a refined Bahadur-Kiefer (BK) expansion. Different from standard BK expansions, we do not bundle together all the higher-order terms (as opposed to, for example, Zhou and Portnoy, 1996; Ota et al., 2019). Notice that the dominant nonlinear term in the BK expansion, $n^{-1/2}G^{-1}B_{n}(\hat{\theta}_{\tau})$, has order $O_{p}\left(n^{-3/4}\sqrt{\log n}\right)$ (Equation (6)). Theorem 3 in Knight (2002) shows that for classical QR with discrete covariates, this term converges in distribution to a zero mean random process. Therefore, we explicitly extract the higher-order terms up to order $O_{p}\left(n^{-1}\right)$ (ignoring logarithmic terms) from the BK remainder. As we will show in the following sections, these higher-order terms admit feasible counterparts.
Remark 1 (Alternative approach for deriving stochastic expansions).
Portnoy (2012) proposed an alternative approach for deriving a stochastic expansion of classical QR estimators. This approach yields bounds on the precision of a nonlinear Gaussian approximation of order $O_{p}\left(n^{-1}\log^{5/2}n\right)$. As we will show, the expansion in Theorem 1 yields a bias formula for both QR and IVQR estimators that admits a feasible implementation. The results in Portnoy (2012) are specific to classical QR, and it is not clear to us whether these results can be used for bias correction using a Nagar-style approach. ∎
Remark 2 (General IVQR estimators).
While we focus on exact IVQR estimators in the main text, the results in Theorem 1 can be used to obtain a uniform BK expansion for general 1-step corrected IVQR estimators. See Appendix E for details. ∎
3.2 Bias formula for exact estimators
Following common practice (e.g., Nagar, 1959; Kaplan and Sun, 2017), for a generic estimator $\hat{\gamma}$, we define the second-order bias $\operatorname{Bias}(\hat{\gamma})$ as the bias of the leading terms in the stochastic expansion of $\hat{\gamma}$ up to the order $O_{p}\left(n^{-1}\right)$. This second-order bias can be interpreted as an approximation of the actual bias that works with arbitrarily high probability in large samples.
Before stating the result, we observe that under our Assumption 2, the moment condition (1) is equivalent to
$\mathbb{E}[(1\{-Y\leq W^{\prime}(-\theta_{\tau})\}-(1-\tau))Z]=0.$
Thus, we can characterize the QR and IVQR estimators using the moment function $g_{\tau}^{*}(\theta)\triangleq\mathbb{E}[(1\{-Y\leq W^{\prime}\theta\}-(1-\tau))Z]$ with sample analog $\hat{g}_{\tau}^{\ast}(\theta)$. The following theorem characterizes the second-order bias in terms of $\hat{g}_{\tau}(\hat{\theta})$ and $\hat{g}_{\tau}^{*}(-\hat{\theta})$. As in Section 3.1, we define $\partial_{\theta}G_{j}(\theta)\triangleq\partial_{\theta}\partial_{\theta}g_{j}(\theta)$ for all $\theta\in\Theta$.
Theorem 2.
Suppose that Assumptions 13 hold. Consider $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ obtained from program (3) for some $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,QR}$. Then the second-order bias is
$\operatorname{Bias}(\hat{\theta}_{\tau})=G^{-1}(\theta_{\tau})\left[\frac{1}{2}\mathbb{E}\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)-\frac{\kappa_{\tau}}{n}-\frac{1}{2n}Q^{\prime}vec(\Omega_{\tau})\right],$ (8)
where
$\kappa_{\tau}$ $\triangleq\left(\tau-\frac{1}{2}\right)\mathbb{E}f_{\varepsilon_{\tau}}(0|W,Z)ZW^{\prime}G^{-1}Z,$
$\Omega_{\tau}$ $\triangleq\operatorname{Var}[Z(1\{Y\leq W^{\prime}\theta_{\tau}\}-\tau)],$
[…]
The second-order bias formula (8) has three components.
The first component, $G^{-1}\mathbb{E}\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)/2$, captures the bias from the sample moments not being zero at the estimator. This term is not equal to zero in general. By Theorem 1 and Assumption 3, this term has bias of order at most $O\left(n^{-1+2/\gamma}\log n\right)$.
The second component, $n^{-1}G^{-1}\kappa_{\tau}$, appears because of the discontinuity in the sample moment functions. The term $\kappa_{\tau}$ reflects the dependence between the sample moments and the linear influence of a single observation on $\hat{\theta}_{\tau}$. Notice that the first two components combined correspond to the second-order bias of the terms $G^{-1}(\theta_{\tau})(\hat{g}_{\tau}(\hat{\theta}_{\tau})-n^{-1/2}B_{n}(\hat{\theta}_{\tau}))$ in Theorem 1.
The last component, $(2n)^{-1}G^{-1}Q^{\prime}vec(\Omega)$, stems from the non-uniformity of the conditional distribution of $Y$ given $(W,Z)$. This term corresponds to the term $G^{-1}(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})/2$ in Theorem 1. Similar terms are typically present in most nonlinear estimators with nonzero Hessian of the score function (see, for example, Rilstone et al., 1996).
To illustrate the approximate bias formula, consider an order statistic of $Y\sim\text{Uniform}(0,1)$ (corresponding to our framework with $Z=W=1$), for which an exact bias formula is available (e.g., Ahsanullah et al., 2013). We show in Appendix C that the precision of the second-order bias formula in this case is $O\left(n^{-2}\right)$, which is smaller than the order of the remainder term in the stochastic expansion of Theorem 1. Figure 1 illustrates the precision of the asymptotic formula by comparing it to the actual bias of the order statistic.
Figure 1: Comparison to exact formula in univariate case
Notes: Exact (circles) and second-order (crosses) biases, scaled by $n$, as functions of quantile level $\tau$ for $\hat{\theta}_{\tau}=Y_{(\lfloor\tau n\rfloor)}$, where $Y\sim\text{Uniform}(0,1)$, $n=10$.
It is interesting to compare our results to the higher-order bias analysis of non-smooth estimators based on the generalized functions heuristic (e.g., Phillips, 1991). In recent work, Lee et al. (2017, 2018) derived a second-order bias formula for classical QR and IVQR under the assumption that the sample moments are zero at the estimator so that the first term of the bias formula vanishes. We show in Appendix C that this term is non-negligible even in simple cases (see also Figure 6 in Section 5).
3.3 Feasible bias correction
The bias formula suggests the following feasible bias-corrected estimator,
$\hat{\theta}_{bc}=\hat{\theta}_{\tau}-\frac{1}{2}\hat{G}^{-1}\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right]+\frac{1}{n}\hat{G}^{-1}\left[\hat{\kappa}_{\tau}+\frac{1}{2}\hat{Q}^{\prime}vec(\hat{\Omega})\right],$
where $\hat{G}$, $\hat{\kappa}_{\tau}$, $\hat{Q}$, and $\hat{\Omega}$ are estimators of $G$, $\kappa_{\tau}$, $Q$, and $\Omega$, respectively, satisfying the following consistency requirement.
Assumption 4 (Consistency of component estimators).
The estimators $\hat{G}$, $\hat{\kappa}_{\tau}$, $\hat{Q}$, and $\hat{\Omega}$ are consistent for $G$, $\kappa_{\tau}$, $Q$, and $\Omega$, respectively. Moreover, $\hat{G}-G=o_{p}\left((n^{1/3}\log n)^{-1}\right)$.
In Section 3.4, we propose finite-difference estimators for which Assumption 4 holds under Assumptions 13. Assumption 4 could also be verified for other nonparametric estimators of the bias components.
The next theorem shows that the second-order bias of the bias-corrected estimator is zero.
Theorem 3.
Suppose that Assumptions 14 hold. Consider $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ obtained from program (3) for some $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,QR}$. Then the feasible bias correction eliminates the second-order bias, $\operatorname{Bias}\left(\hat{\theta}_{bc}\right)=0.$
Note that the requirement that $\hat{G}-G=o_{p}\left((n^{1/3}\log n)^{-1}\right)$ in Assumption 4 is necessary to ensure that the contribution of the product of the sample moments, which are $O_{p}\left(n^{-1+2/\gamma}\log n\right)$ with $\gamma\geq 6$ (as required by Assumption 3) by Theorem 1, and the estimation error in $G^{-1}$ can be omitted in computing the second-order bias. This condition is only required for IVQR estimators. For classical QR estimators, the sample moments are of order $O_{p}\left(n^{-1}\right)$, so that consistency of $\hat{G}$ at any rate of convergence suffices for Theorem 3.
3.4 Finite difference estimators of bias components
To implement the bias correction, we need estimators of $G$, $\kappa_{\tau}$, $Q$, and $\Omega$ that satisfy Assumption 4. The variance matrix $\Omega$ can be estimated using the analogy principle,
$\hat{\Omega}_{\tau}\triangleq\mathbb{E}_{n}[Z(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)-\mathbb{E}_{n}Z(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)]^{2}.$
All other bias components take the form of derivatives. Therefore, we leverage our theoretical results on the properties of the sample moments to develop a unified finite-difference framework for estimating these components.
Under Assumptions 1 and 2, the Jacobian is $G=\mathbb{E}f_{\varepsilon_{\tau}}(0|W,Z)ZW^{\prime}$ and the Hessian consists of gradients of the components of $G$, i.e. $\partial_{\theta}G_{i,j}(\theta_{\tau})=\mathbb{E}f^{(1)}_{\varepsilon_{\tau}}(0|W,Z)Z_{i}W_{j}W$, where $i,j=1,\dots,k.$ This suggests the following analog estimators.
The $(i,j)$-th component of $G$ can be estimated using Powell (1986)’s estimator
$\hat{G}_{i,j}=\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{1,n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{1,n}\}}{2h_{1,n}}Z_{i}W_{j}\right],$ (9)
where $h_{1,n}\to 0$ is a bandwidth. Denote by $e_{\ell}$ the $\ell$-th unit vector in $\mathbb{R}^{k}$, where $\ell=1,\dots,k$. The derivative of the $(i,j)$-th component of $G$ in the direction $e_{\ell}$ (i.e. the second partial derivative of $g_{\tau}$) can be estimated as the symmetric first difference of (9),
$\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ $=\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{2,n}\}-2\cdot 1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}+1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{2,n}\}}{h_{2,n}^{2}}Z_{i}W_{j}W_{\ell}\right],$
where $h_{2,n}\to 0$ is a (potentially different) bandwidth. For $\kappa_{\tau}$, the finite difference sample analog is
$\hat{\kappa}_{\tau}=\left(\tau-\frac{1}{2}\right)\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{3,n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{3,n}\}}{2h_{3,n}}ZW^{\prime}\hat{G}^{-1}Z\right],$
where $h_{3,n}\to 0$ is a bandwidth. Finally, $Q$ can be estimated by the sample analog matrix $\hat{Q}$ with columns
$\hat{Q}_{j}\triangleq vec\left[(\hat{G}^{-1})^{\prime}\widehat{\partial_{\theta}G_{j}}\hat{G}^{-1}\right],\quad j=1,\dots,k,$
where $\widehat{\partial_{\theta}G_{j}}$ is the matrix with elements $\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ for $i,\ell=1,\dots,k$.
The next lemma establishes the consistency of the estimators of the bias components. It implies that these estimators satisfy the high-level conditions in Assumption 4. Moreover, it provides nearly remainder-optimal bandwidth rates, i.e., the rates that yield the fastest convergence rates of the stochastic remainder terms of the corresponding stochastic expansions (up to logarithmic terms).
Lemma 1.
Suppose that Assumptions 13 hold. Then the nearly remainder-optimal bandwidth rates are $h_{1,n}\propto n^{-1/5}$, $h_{2,n}\propto n^{-1/7}$, $h_{3,n}\propto n^{-1/5}$, and, under these bandwidth rates,
$\hat{G}$ $=G+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right),$
$\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ $={({\partial_{\theta}G}_{i,j})}_{\ell}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/7}}\right),$
$\hat{Q}_{j}$ $=Q_{j}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/7}}\right),$
$\hat{\kappa}_{\tau}$ $=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right),$
$\hat{\Omega}_{\tau}$ $=\Omega+O_{p}\left(\frac{1}{\sqrt{n}}\right).$
Moreover, the convergence rate for $\hat{G}$ is uniform in $\tau\in[\varepsilon,1-\varepsilon]$.
Proposition 1 in Kato (2012) shows that the remainder rate for $\hat{G}$ in Lemma 1, $h_{1,n}\propto n^{-1/5}$, is the AMSE-optimal rate.666 We conjecture that analogous AMSE-optimality results could be established for estimators $\widehat{({\partial_{\theta}G}_{i,j})}_{\ell}$ and $\hat{Q}_{j}$, but leave this extension for future work.
To implement the finite difference estimators in practice, one needs to choose constants in addition to the bandwidth rates. A natural approach would be to use non-parametric methods such as cross-validation to estimate the optimal bandwidth parameter. However, such non-parametric methods typically require large sample sizes (e.g. Simonoff, 1996), which makes them impractical for our purposes. Another common approach, which we take here, is to propose rule-of-thumb choices that account for the scale of the data. To be robust to heavy-tailed distributions, we suggest using a robust measure of dispersion. Specifically, we propose a rule-of-thumb choice in form $h_{1,n}=A_{G}\cdot n^{-1/5}$, $h_{2,n}=A_{Q}\cdot n^{-1/7}$, $h_{3,n}=A_{\kappa}\cdot n^{-1/5}$, where, for all $j\in\{G,Q,\kappa\}$,
$A_{j}=\tilde{A}_{j}\cdot 1.48\cdot\widehat{\operatorname{MAD}}_{\tau},$
and $\widehat{\operatorname{MAD}}_{\tau}$ is the estimated median absolute deviation of the $\tau$-quantile residuals. The constant $1.48$ is chosen because for the normal distribution 1.48 times the median absolute deviation is equal to the standard deviation. [^7] Chernozhukov et al. (2013) suggest using $\operatorname{IQR}/1.35$ as a robust estimate of the scale parameter. For the normal distribution, this choice coincides with $1.48\operatorname{MAD}$. We suggest choosing $\tilde{A}_{G}=\tilde{A}_{\kappa}=2$ and $\tilde{A}_{Q}=1.5$ based on our simulation evidence, as we describe in more detail in Section 4.2.
4 Simulation evidence
In this section, we evaluate the performance of our feasible bias correction procedure in a Monte Carlo simulation study.
4.1 Simulation design
We consider data-generating processes (DGPs) inspired by the simulations in Andrews and Mikusheva (2016). The outcome is generated according to the following location-scale model
$Y_{i}=W_{i}+(0.5+W_{i})U_{i},\quad i=1,\dots,n,$
where $W_{i}=\Phi(\tilde{W}_{i})$, $Z_{i}=\Phi(\tilde{Z}_{i})$, $U_{i}=F_{U}^{-1}\left(\Phi(\tilde{U}_{i})\right)$, $(\tilde{W}_{i},\tilde{Z}_{i},\tilde{U}_{i})\sim N(0,\Sigma)$, $\Sigma_{11}=\Sigma_{22}=\Sigma_{33}=1$, $\Sigma_{23}=0$, and $\Phi$ is the standard normal CDF. Hence, in all the designs, both the regressors and the instruments are $\operatorname{Uniform}[0,1]$. We consider six DGPs that differ with respect to the error distribution $F_{U}$ and whether or not $W$ is exogenous.
[…]
In Appendix F, we consider two additional DGPs to assess the impact of the strength of the instrument.
4.2 Bandwidth choice
An important practical issue is the choice of the bandwidths $h_{1,n}$, $h_{2,n}$, and $h_{3,n}$. We use the (data-dependent) rule-of-thumb bandwidth choice described in Section 3.4. To implement this rule-of-thumb bandwidth choice, we need to choose the constants $\tilde{A}_{j}$, $j\in\{G,Q,\kappa\}$. We found that across the DGPs considered in our simulations, $\tilde{A}_{G}=\tilde{A}_{\kappa}=2$ and $\tilde{A}_{Q}=1.5$ perform well in terms of bias across all designs, especially when the bias is large. [^8] If more information about the DGP is available, it is possible to further refine the tuning parameter. For example, under DGP1 where $Q=0$, choosing smaller values for $A_{Q}$ results in a better alignment with the infeasible formula. However, absent such information, we consider the proposed rule of thumb to be a reasonable compromise when one cares about the performance across different designs, especially those with larger biases (e.g., DGP3–DGP6).
We compare the performance of the feasible bias correction based on the rule-of-thumb bandwidth choice to the performance of the corresponding infeasible bias correction based on the true $G$, $Q$, and $\kappa$. For IVQR, the exact formulas are not available. Therefore, we use numerically computed values based on 10 million observations and tuning choice ${A}_{G}={A}_{Q}={A}_{\kappa}=1$.
4.3 Results
We focus on the performance of bias correction for $\tau\in\{0.25,0.5,0.75\}$. Figure 2 shows the impact of bias correction for the exogenous DGP1–DGP3 with $n=100$. Figure 3 shows the corresponding results for the endogenous DGPs (DGP4–DGP6). We use classical QR of $Y$ on $W$, implemented via the linear programming formulation in Appendix D, for DGP1–DGP3 and IVQR, implemented using the MILP formulation in Appendix D, for DGP4–DGP6.
The main findings can be summarized as follows. First, the QR estimators based on the exogenous designs (DGP1–DGP3) exhibit lower biases than the IVQR estimators based on the corresponding endogenous designs (DGP4–DGP6). For both estimators, the biases tend to be the largest for the designs with heavy tails (DGP3 and DGP6). Second, the feasible bias correction based on the rule-of-thumb bandwidth choice reduces the bias of QR and IVQR estimators in many cases. The bias reductions are the most notable when the biases of the original estimators are large, which is when bias correction is most needed. Finally, the infeasible bias correction reduces the bias in most cases, underscoring the usefulness of the proposed theory. See Appendix F for additional simulation evidence on the impact of the sample size and instrument strength on the performance of bias correction.
Next, we investigate the impact of bias correction on the RMSE of the estimators. The proposed bias correction approach is designed to reduce the bias but is not theoretically guaranteed to reduce the RMSE. Figure 4 reports the results for DGP1 and DGP4 with $n=100$. Overall, the impact of bias correction on the RMSE is rather small. While the infeasible bias correction can slightly increase the RMSE for DGP1, it decreases the RMSE across all quantile levels for DGP4. The feasible bias correction with the rule-of-thumb bandwidth slightly increases the RMSE in most cases. Appendix Figures 10 and 11 show the results for the other DGPs, and Appendix Figure 12 shows the corresponding results for the Mean Absolute Deviation, an alternative measure of risk.
Finally, we study the impact of the bias correction on the coverage probability of the standard confidence intervals. To focus on the impact of bias correction, we use the same standard errors based on the rule-of-thumb bandwidth for the original and the bias-corrected estimator. As a result, by construction, the bias correction does not affect the length of the confidence intervals. Figure 5 shows the empirical coverage for DGP1 and DGP4 before and after bias correction. The feasible bias correction based on the rule-of-thumb bandwidth leads to higher coverage accuracy in the majority of cases but can lead to some undercoverage at the median. Thus, while bias correction may not improve the RMSE in small samples, it can lead to more accurate inferences. The results for the other DGPs are in Appendix Figures 13 and 14, and Appendix Figure 15 presents the results for $n=200$.
Figure 2: Bias (multiplied by $n$) before and after correction for DGP1–DGP3
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and QR with infeasible bias correction (gold dashed line) for DGP1–DGP3. All results are based on 5,000 simulation repetitions.
Figure 3: Bias (multiplied by $n$) before and after correction for DGP4–DGP6
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP4–DGP6. All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
Figure 4: RMSE comparison of raw and bias-corrected estimators
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP1, classical QR and (b) DGP4, IVQR. All results are based on 5,000 simulation repetitions.
Figure 5: Confidence interval coverage before and after correction for DGP1 and DGP4
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP1 (classical QR) and DGP4 (IVQR). All results are based on 5,000 simulation repetitions.
5 Empirical application
The second-order bias matters most in applications with small sample sizes. We therefore illustrate our bias correction approach using the classical dataset of Engel (1857), analyzed by Koenker and Bassett (1982) and Koenker and Hallock (2001), among others. The data contain information on annual income and food expenditure (in Belgian francs) for $n=235$ Belgian working-class households and are obtained from the R package quantreg (Koenker, 2022). One feature of these data is the growing dispersion of the outcome variable (food expenditure) as a function of the regressor (income) (Koenker and Hallock, 2001), which is similar to our Monte Carlo designs. We divide the values of income and food expenditure by $1000$ so that the unit of measurement becomes a thousand Belgian francs. This makes the scale of intercept and slope parameters comparable.
Figure 6: Quantile regression of annual food expenditure on income
(a) Impact of bias correction
(b) Composition of second-order bias
Notes: Panel (a) compares the classical QR estimates (blue dots) to the bias-corrected QR estimates (gold squares) with 90% confidence intervals (bars). The bias correction is based on the rule-of-thumb bandwidth choice with $(\tilde{A}_{G},\tilde{A}_{Q},\tilde{A}_{\kappa})=(2,1.5,2)$. Panel (b) shows the contributions of the different bias components to the overall second-order bias.
We estimate classical QRs of food expenditure ($Y$) on income ($W$) and a constant (blue dots). The bias-corrected estimates (gold squares) are obtained using the recommended rule-of-thumb bandwidth choice th $(\tilde{A}_{G},\tilde{A}_{Q},\tilde{A}_{\kappa})=(2,1.5,2)$. Figure 6 presents the results. Panel (a) compares the classical and the bias-corrected QR estimates. The results suggest that the impact of bias correction is more pronounced around the median and in the tails. The magnitude of the differences between the original and bias-corrected estimates can be larger than 50% of the standard errors, which is substantial given that we are focusing on the second-order bias.
Panel (b) shows the individual contributions of the different components to the overall second-order bias. We can decompose the bias correction term as follows:
$\hat{\theta}_{\tau}-\hat{\theta}_{bc}=\underbrace{\frac{1}{2}\hat{G}^{-1}\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right]}_{\text{(i)}}-\underbrace{\frac{1}{n}\hat{G}^{-1}\hat{\kappa}_{\tau}}_{\text{(ii)}}-\underbrace{\frac{1}{2n}G^{-1}\hat{Q}^{\prime}vec(\hat{\Omega})}_{\text{(iii)}}.$
The main takeaway from the bias decomposition is that while all three components play a role, the sample moment term (i) and especially the Hessian term (iii) can be large and account for most of the bias. The $\kappa$-term (ii) is smaller overall and only matters in the tails.
6 Conclusion
We demonstrate that classical QR and IVQR estimators can exhibit a non-negligible second-order bias. We characterize this bias theoretically and use this characterization to derive a novel analytical bias correction method. The proposed feasible bias correction reduces the bias of QR and IVQR estimators across a variety of settings at a very low computational cost. However, there is scope for further improving its performance when the sample size is very small and the instruments are weak. The simulation performance of the infeasible bias correction based on the population version of our theoretical bias formula suggests that exploring alternative feasible bias correction approaches is a promising direction for future research. For example, one could consider regularization approaches or explore imposing parametric assumptions to improve the estimation of the bias components.
Acknowledgments
We are grateful to the Editor (Xiaohong Chen), the Associate Editor, and three anonymous referees, as well as Victor Chernozhukov, Zheng Fang, Dalia Ghanem, Jiaying Gu, Marc Henry, Keisuke Hirano, Nail Kashaev, Roger Koenker, Vladimir Koltchinskii, Michal Kolesar, Simon Lee, Blaise Melly, Hyungsik Roger Moon, Hashem Pesaran, Joris Pinkse, Wolfgang Polonik, Stephen Portnoy, Geert Ridder, Andres Santos, Davide Viviano, Yuanyuan Wan, and seminar participants at UC Berkeley, UC Davis, University of Toronto, and USC for valuable comments. All errors and omissions are our own.
Appendix A Bahadur-Kiefer representation, proofs
A.1 Auxiliary results for generic IVQR estimators
Lemma A.1.
Under Assumptions 2 and 3, $g_{\tau}(\theta)$ is three times continuously differentiable in $\theta$.
Proof.
By definition, $g_{\tau}(\theta)=\mathbb{E}(1\{Y\leq W^{\prime}\theta\}-\tau)Z=\mathbb{E}(\mathbb{E}(F_{Y}(W^{\prime}\theta|W,Z)-\tau)Z)$. The result then follows from the dominated convergence theorem. ∎
Lemma A.2.
Suppose Assumptions 2 and 3 hold. Then for any estimator $\hat{\theta}_{\tau}$ such that $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|=O_{p}\left({r^{-1}_{n}}\right)$ for some sequence $r_{n}\to\infty$, we have a representation
$\hat{g}_{\tau}(\hat{\theta}_{\tau})$ $=\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})+\tau(\mathbb{E}Z-\mathbb{E}_{n}{Z})+\frac{1}{\sqrt{n}}B_{n}(\hat{\theta}_{\tau})$
$+G(\theta_{\tau})(\hat{\theta}_{\tau}-\theta_{\tau})+\frac{1}{2}(\hat{\theta}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau}-\theta_{\tau})+O_{p}\left(\frac{1}{r^{3}_{n}}\right),$ (10)
where the remainder rate is uniform in $\tau\in[\varepsilon,1-\varepsilon]$.
Proof.
By definition,
$\hat{g}_{\tau}(\hat{\theta}_{\tau})$ $=\mathbb{E}_{n}1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}Z-\tau\mathbb{E}_{n}Z$
$=\frac{1}{\sqrt{n}}B^{\circ}_{n}(\hat{\theta}_{\tau})+g^{\circ}(\hat{\theta}_{\tau})-\tau\mathbb{E}_{n}Z$
$=\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})+\frac{1}{\sqrt{n}}B_{n}(\hat{\theta}_{\tau})+\tau(\mathbb{E}Z-\mathbb{E}_{n}{Z})+g_{\tau}(\hat{\theta}_{\tau}).$
By Lemma A.1, $g_{\tau}(\cdot)$ is three times continuously differentiable. Since $\theta$ is restricted to a compact set $\Theta$, the norm of the third derivative is bounded on $\Theta$. The Taylor theorem implies that there exist a neighborhood of $\theta_{\tau}$ such that for any $\theta$ in the neighborhood,
$g_{\tau}(\theta)=G(\theta_{\tau})(\theta-\theta_{\tau})+\frac{1}{2}(\theta-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\theta-\theta_{\tau})+R(\theta),$
where $R(\theta)=O\left(\|\theta-\theta_{\tau}\|^{3}\right)$ uniformly in $\tau$. Then (10) follows immediately because $\hat{\theta}_{\tau}$ is a uniformly consistent estimator. ∎
Now let us study the large sample behavior of the term $B_{n}(\hat{\theta}_{\tau})$ in (10).
Lemma A.3.
Suppose that Assumptions 2 and 3 hold. For any pair of estimators $\hat{\theta}_{\tau}$ and $\hat{\theta}_{\tau}^{\ast}$ such that $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}^{\ast}-\hat{\theta}_{\tau}\|=O_{p}\left({r^{-1}_{n}}\right)$ for some sequence $r_{n}\to\infty$, we have
$B_{n}(\hat{\theta}_{\tau})-B_{n}(\hat{\theta}_{\tau}^{\ast})=O_{p}\left(\sqrt{\frac{{\log r_{n}}}{r_{n}}}\right)+o_{p}\left(\frac{\log r_{n}}{n^{1/3}}\right)\quad\text{uniformly in }\tau\in[\varepsilon,1-\varepsilon].$
Proof.
The proof relies on the arguments in Ota et al. (2019, Lemma 3) adapted to our setting. The idea is to verify the conditions of Lemma 1 of Ota et al. (2019), which follows from Corollary 5.1 in Chernozhukov et al. (2014) and use this corollary to prove the desired result.
Since $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}^{\ast}-\hat{\theta}_{\tau}\|=O_{p}\left({r^{-1}_{n}}\right)$, we have $P\left(\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}^{\ast}-\hat{\theta}_{\tau}\|\leq{M_{n}}/{r_{n}}\right)\to 1$ for any sequence $M_{n}\to\infty$. Consider the functions
$f_{\theta,h,\alpha}:\,(y,w,z)\mapsto\left(1\{y-w^{\prime}\theta\leq w^{\prime}h\}-1\{y-w^{\prime}\theta\leq 0\}\right)\alpha^{\prime}z$
that constitute the function class
$\mathcal{F}_{n}$ =$=\bigg{\{}f_{\theta,h,\alpha}:\,\,\theta\in\Theta,\|h\|\leq\frac{M_{n}}{r_{n}},\|\alpha\|=1\bigg{\}},$
where $M_{n}$ is a sequence such that $M_{n}\to\infty$ and $M_{n}/r_{n}\to 0$. Let $\mathbb{G}_{n}$ be the standard empirical process operator on $\mathcal{F}_{n}$ with the data $(Y_{i},W_{i},Z_{i})$, $i=1,\dots,n$. By Assumption 3, $\mathcal{F}_{n}$ admits an envelope $F(y,w,z)\equiv\|z\|$.
Let us now verify the conditions in Lemma 1 of Ota et al. (2019). First, since
$\left(1\left\{Y-W^{\prime}\theta\leq W^{\prime}h\right\}-1\left\{Y-W^{\prime}\theta\leq 0\right\}\right)^{2}=1\left\{\min(0,W^{\prime}h)<Y-W^{\prime}\theta\leq\max(0,W^{\prime}h)\right\},$
we obtain
$\mathbb{E}f_{\theta,h,\alpha}^{2}(Y,W,Z)$ $=\mathbb{E}\left[\left(1\big{\{}Y-W^{\prime}\theta\leq W^{\prime}h\big{\}}-1\big{\{}Y-W^{\prime}\theta\leq 0\big{\}}\right)\alpha^{\prime}Z\right]^{2}$
$\leq\mathbb{E}\|Z\|^{2}\cdot\mathbb{E}1\bigg{\{}\min(0,W^{\prime}h)<Y-W^{\prime}\theta\leq\max(0,W^{\prime}h)\bigg{\}}$
$\leq km^{2/\gamma}\cdot\mathbb{E}1\bigg{\{}\min(0,W^{\prime}h)<Y-W^{\prime}\theta\leq\max(0,W^{\prime}h)\bigg{\}}$
$=km^{2/\gamma}\cdot\mathbb{E}\bigg{(}\mathbb{E}\left[\left|F_{Y}(W^{\prime}h+W^{\prime}\theta|W,Z)-F_{Y}(W^{\prime}\theta|W,Z)\right|\,\bigg{|}\,W,Z\right]\bigg{)}$
$\leq km^{2/\gamma}\cdot\mathbb{E}\bigg{(}\mathbb{E}\left[|W^{\prime}h|\sup_{y}f_{Y}(y|W,Z)\,\bigg{|}\,W,Z\right]\bigg{)}\leq k^{2}m^{4/\gamma}\bar{f}\|h\|=O\left(\frac{M_{n}}{r_{n}}\right),$
where we used Assumptions 2 and the inequality
$\mathbb{E}\|Z\|^{2}\leq k\max_{j=1,\dots,k}\mathbb{E}|Z_{j}|^{2}\leq k\max_{j=1,\dots,k}\left(\mathbb{E}|Z_{j}|^{\gamma}\right)^{2/\gamma}$
in conjunction with Assumption 3.
Therefore, the variance parameter of the process is
$\sigma_{n}^{2}\triangleq\sup_{f\in\mathcal{F}_{n}}\mathbb{E}f^{2}(Y,W,Z)=O\left(\frac{M_{n}}{r_{n}}\right).$
Second, using Lemma 2 in Ota et al. (2019), we have
$\mathbb{E}\max_{1\leq i\leq n}F^{2}(Y_{i},W_{i},Z_{i})=\mathbb{E}\max_{1\leq i\leq n}\|Z_{i}\|^{2}=o(n^{2/\gamma}).$
Third, because the function class $\mathcal{F}_{n}$ is a VC class (Vapnik and Chervonenkis, 1971) with the envelope $\|Z\|$, there exist constants $A$ and $V$ independent of $n$ such that the standard entropy bound
$\sup_{Q}N\left(\mathcal{F}_{n},\|\cdot\|_{Q,2},\eta\|Z\|_{Q,2}\right)\leq(A/\eta)^{V},\text{ for all }\eta\in(0,1]$
holds (e.g., van der Vaart and Wellner, 1996, Section 2.6). Here the supremum is taken over all finitely discrete measures $Q$ and $\|\cdot\|_{Q,2}$ is the $L^{2}(Q)$ norm.
Finally, applying Lemma 1 in Ota et al. (2019), we obtain
$\mathbb{E}\sup_{\theta\in\Theta,\|h\|\leq\frac{M_{n}}{r_{n}},\|\alpha\|=1}\|\mathbb{G}_{n}f_{\theta,h,\alpha}\|\lesssim\sqrt{V\sigma_{n}^{2}\log(Am/\sigma_{n})}+\frac{V\sqrt{\mathbb{E}\max_{1\leq i\leq n}\|Z_{i}\|^{2}}}{\sqrt{n}}\log(Am/\sigma_{n})$
$=O\left(\sqrt{\frac{\log r_{n}}{r_{n}}}\right)+o\left(\frac{\log r_{n}}{n^{\frac{\gamma-2}{2\gamma}}}\right),$ (11)
where the last equality holds by choosing $M_{n}\to\infty$ sufficiently slowly. Note that the right-hand side of this equation does not depend on $\tau$. Consequently, by the definition of the norm and equation (11),
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|B_{n}(\hat{\theta}_{\tau})-B_{n}(\hat{\theta}_{\tau}^{*})\|$ $=\sup_{\tau\in[\varepsilon,1-\varepsilon]}\max_{\|\alpha\|=1}\bigg{|}\mathbb{G}_{n}f_{\hat{\theta}_{\tau},(\hat{\theta}_{\tau}-\hat{\theta}_{\tau}^{*}),\alpha}\bigg{|}$
$=O_{p}\left(\sqrt{\frac{{\log r_{n}}}{r_{n}}}\right)+o_{p}\left(\frac{\log r_{n}}{n^{\frac{\gamma-2}{2\gamma}}}\right),$
where the last equality holds by Markov’s inequality.
A.2 Auxiliary results for exact estimators
Lemma A.4.
Under Assumptions 1.1, 2 and 3, $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|=o_{p}\left(1\right)$, where $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ for any $p\in[1,\infty]$ or $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,\text{QR}}$.
Proof.
We give the proof for $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$. Uniform consistency for the case of $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,\text{QR}}$ was established by Angrist et al. (2006, Theorem 3).
By Assumption 1.1,
$\arg\min_{\theta\in\Theta}\|g_{\tau}(\theta)\|_{p}=\theta_{\tau}.$
Assumptions 2 and 3 imply that the function class
$\left\{(w,y,z)\mapsto z(1\{y\leq w^{\prime}\theta\}-\tau),\leavevmode\nobreak\ \tau\in[\varepsilon,1-\varepsilon],\leavevmode\nobreak\ \theta\in\Theta\right\}$
is Donsker (compare with the function class $\mathcal{F}_{n}$ in the proof of Lemma A.3) and thus Glivenko-Cantelli, and hence
$\sup_{\theta\in\Theta,\tau\in[\varepsilon,1-\varepsilon]}|\hat{g}_{\tau}(\theta)-g_{\tau}(\theta)|=\sup_{\theta\in\Theta,\tau\in[\varepsilon,1-\varepsilon]}|(\mathbb{E}_{n}-\mathbb{E})Z_{i}(1\{Y_{i}\leq W^{\prime}\theta\}-\tau)|\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}0.$
By the argmin theorem (e.g., Theorem 2.1 in Newey and McFadden, 1994) applied to $Q_{n}(\theta,\tau)\triangleq\|\hat{g}_{\tau}(\theta)\|_{p}$, we get $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau,p}-\theta_{\tau}\|=o_{p}\left(1\right)$. ∎
Lemma A.5.
Under Assumptions 13, for any exact QR estimator $\hat{\theta}_{\tau,\text{QR}}$ as defined in equation (2), we have
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau,\text{QR}}-\theta_{\tau}\|$ $=O_{p}\left(\frac{1}{\sqrt{n}}\right),$ (12)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,\text{QR}})\|_{p}$ $=O_{p}\left(\frac{k}{n}\right).$
Proof.
Equation (12) follows from Theorem 3 in Angrist et al. (2006).
The exact QR estimators yield exact zeros of the subgradient
$\frac{1}{n}\sum_{i=1}^{n}(\tau-h(Y_{i}-W_{i}^{\prime}\theta))W_{i},$
where the multi-valued function $h(u)$ is defined as $1\{u<0\}$ for $u\neq 0$ and $h(0)\triangleq[0,1]$ for $u=0$. The subgradient function differs from sample moment functions by the fraction of observations with $Y_{i}=W_{i}^{\prime}\hat{\theta}_{\tau,\text{QR}}$. Note that under Assumption 2, $Y_{i}$ has density with respect to the Lebesgue measure conditional on $W_{i}$. Thus the observations $(Y_{i},W_{i})$ are in general position with probability 1 (see Definition 2.1 and the subsequent discussion in Koenker, 2005). Because the observations are in general position with probability 1, there are at most $k$ terms like that, and so $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,\text{QR}})\|_{p}=O_{p}\left({k}/{n}\right)$. ∎
Lemma A.6.
Under Assumptions 1, 2, and 3, for any estimator $\hat{\theta}_{\tau}=\hat{\theta}_{\tau,p}$ that minimizes $\|\hat{g}_{\tau}(\theta)\|_{p}$, we have
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|$ $=O_{p}\left(\frac{1}{\sqrt{n}}\right),$ (13)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\hat{\xi}_{\tau}\|$ $=o_{p}\left(\frac{1}{\sqrt{n}}\right),$ (14)
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau})\|_{p}$ $=o_{p}\left(\frac{1}{\sqrt{n}}\right),$ (15)
where $\hat{\xi}_{\tau}$ is defined in equation (4).
Proof.
The proof proceeds in four steps.
Step 1. Under the assumptions of the lemma, the empirical process $B_{n}^{\circ}(\theta)$ is Donsker (see proof of Lemma A.3 above) and thus asymptotically stochastically equicontinuous (see discussion in van der Vaart and Wellner, 1996, Section 2.1.2; also Theorem 1.5.7 and Problem 2.1.5 in the same book).
Step 2. By definition, $\sqrt{n}(\hat{\xi}_{\tau}-\theta_{\tau})$ can be written as
$\sqrt{n}(\hat{\xi}_{\tau}-\theta_{\tau})$ $=-G^{-1}_{\tau}\sqrt{n}\left[\tau(\mathbb{E}Z-\mathbb{E}_{n}Z)+\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})\right],$ (16)
where $\theta_{\tau}$ and $G_{\tau}\triangleq\partial_{\theta}g_{\tau}(\theta_{\tau})$ are well-defined by Assumption 1 . This class of functions indexed by $\tau$ is Donsker by similar arguments as in Lemma A.3 (note that the parameter $\tau$ enters this class only through $\theta_{\tau}$ and through the linear term $\tau Z_{i}$). The Donsker property implies
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\xi}_{\tau}-\theta_{\tau}\|=O_{p}\left(\frac{1}{\sqrt{n}}\right).$ (17)
By Lemma A.2 applied to $\hat{\xi}_{\tau}$,
$\hat{g}_{\tau}(\hat{\xi}_{\tau})$ $=\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})+\tau(\mathbb{E}Z-\mathbb{E}_{n}{Z})+\frac{1}{\sqrt{n}}B_{n}(\hat{\xi}_{\tau})$
$+g(\theta_{\tau})+G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})+\frac{1}{2}(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})+O_{p}\left(n^{-\frac{3}{2}}\right).$
Then after substituting the first equation in (17) into the term $G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})$, we have
$\hat{g}_{\tau}(\hat{\xi}_{\tau})=\frac{1}{\sqrt{n}}B_{n}(\hat{\xi}_{\tau})+O_{p}\left(\frac{1}{n}\right).$
So by Step 1, $\hat{g}_{\tau}(\hat{\xi}_{\tau})=O_{p}\left(n^{-\frac{1}{2}}\right)$.
Since $\hat{\theta}_{\tau,p}$ is defined as the estimator attaining the minimal $p$-norm,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,p})\|_{p}\leq\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\xi}_{\tau})\|_{p}=O_{p}\left(n^{-\frac{1}{2}}\right).$
Step 3. Consider $\hat{\xi}^{(2)}_{\tau}\triangleq\hat{\xi}_{\tau}-G^{-1}{B_{n}(\hat{\xi}_{\tau})}/{\sqrt{n}}$. By equation (17), $\hat{\xi}_{\tau}$ is uniformly consistent, and hence $\hat{\xi}^{(2)}_{\tau}=\hat{\xi}_{\tau}+o_{p}\left({1}/{\sqrt{n}}\right)$ uniformly in $\tau\in[\varepsilon,1-\varepsilon]$ since $B_{n}$ is stochastically equicontinuous (Step 1). Then by the stochastic equicontinuity of $B^{\circ}_{n}$ (Step 1) and Lemma A.2 applied to $\hat{\xi}^{(2)}_{\tau}$, uniformly in $\tau\in[\varepsilon,1-\varepsilon]$,
$\hat{g}_{\tau}(\hat{\xi}^{(2)}_{\tau})=\frac{B_{n}^{\circ}(\hat{\xi}^{(2)}_{\tau})-B^{\circ}_{n}(\hat{\xi}_{\tau})}{\sqrt{n}}+o_{p}\left(\frac{1}{\sqrt{n}}\right)=o_{p}\left(\frac{1}{\sqrt{n}}\right).$
This implies
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,p})\|_{p}\leq\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\xi}^{(2)}_{\tau})\|_{p}=o_{p}\left(n^{-\frac{1}{2}}\right).$
which establishes (15).
Step 4. By Lemma A.4, $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau,p}-\theta_{\tau}\|=O_{p}\left({r^{-1}_{n}}\right)$ for some $r_{n}\to\infty$. By Lemma A.2 and Steps 1 and 2, $\hat{\theta}_{\tau,p}$ satisfies
$G(\theta_{\tau})(\hat{\theta}_{\tau,p}-\theta_{\tau})+\frac{1}{2}(\hat{\theta}_{\tau,p}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau,p}-\theta_{\tau})$
$=\hat{g}_{\tau}(\hat{\theta}_{\tau,p})-\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})-\tau(\mathbb{E}Z-\mathbb{E}_{n}{Z})-\frac{1}{\sqrt{n}}B_{n}(\hat{\theta}_{\tau,p})+O_{p}\left(\frac{1}{r_{n}^{3}}\right)$
$=O_{p}\left(\frac{1}{\sqrt{n}}\right)+O_{p}\left(\frac{1}{r_{n}^{3}}\right),$ (18)
uniformly in $\tau\in[\varepsilon,1-\varepsilon]$.
By Assumption 1.2, we can multiply the last equation by $G^{-1}(\theta_{\tau})$ and obtain
$\hat{\theta}_{\tau,p}-\theta_{\tau}+O_{p}\left(\frac{1}{r_{n}^{2}}\right)$ $=O_{p}\left(\frac{1}{\sqrt{n}}\right)+O_{p}\left(\frac{1}{r_{n}^{3}}\right),$
which implies we can take $r_{n}=\sqrt{n}$ by a fixed point argument, which is discussed in detail in Step 3 of the proof of Lemma A.7 below. This implies (13).
By uniform consistency of $\hat{\theta}_{\tau,p}$ and Step 1,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|B_{n}(\hat{\theta}_{\tau,p})\|=\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|B_{n}^{\circ}(\hat{\theta}_{\tau,p})-B_{n}^{\circ}(\theta_{\tau})\|=o_{p}(1).$
Lemma A.2 applied to $\hat{\theta}_{\tau,p}$ yields
$\hat{\theta}_{\tau,p}$ $=\hat{\xi}_{\tau}-G^{-1}(\theta_{\tau})\frac{1}{2}(\hat{\theta}_{\tau,p}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau,p}-\theta_{\tau})+G^{-1}(\theta_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau,p})$
$-\frac{1}{\sqrt{n}}G^{-1}(\theta_{\tau})B_{n}(\hat{\theta}_{\tau,p})+O_{p}\left(n^{-3/2}\right).$
The term $G^{-1}(\theta_{\tau})2^{-1}(\hat{\theta}_{\tau,p}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau,p}-\theta_{\tau})$ is $O_{p}(n^{-1})$. The term $G^{-1}(\theta_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau,p})$ is $o_{p}(n^{-1/2})$ by Step 3. The term $n^{-1/2}G^{-1}(\theta_{\tau})B_{n}(\hat{\theta}_{\tau,p})$ is $o_{p}(n^{-1/2})$ by stochastic equicontinuity of $B_{n}$ (Step 1) and uniform consistency of $\hat{\theta}_{\tau,p}$. Therefore,
$\hat{\theta}_{\tau,p}$ $=\hat{\xi}_{\tau}+o_{p}\left(n^{-\frac{1}{2}}\right)$
uniformly in $\tau\in[\varepsilon,1-\varepsilon]$. This proves (14). ∎
The results of the previous lemma can be further refined.
Lemma A.7.
Under Assumptions 13, for any estimator $\hat{\theta}_{\tau,p}$ that minimizes $\|\hat{g}_{\tau}(\theta)\|_{p}$, we have
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau,p})\|_{p}$ $=O_{p}\left(\frac{\log n}{n^{1-\frac{2}{\gamma}}}\right).$
Proof.
The proof proceeds in four steps.
Step 1. By Lemma A.3 and A.6 applied to $\hat{\theta}_{\tau,p}$ and $\theta_{\tau}$,
$B_{n}(\hat{\theta}_{\tau,p})=B_{n}(\hat{\theta}_{\tau,p})-B_{n}(\theta_{\tau})=O_{p}\left(\sqrt{\frac{{\log\sqrt{n}}}{\sqrt{n}}}\right)+o_{p}\left(\frac{\log n}{n^{\frac{\gamma-2}{2\gamma}}}\right)\quad\text{uniformly in }\tau\in[\varepsilon,1-\varepsilon].$
Since $\frac{\gamma-2}{2\gamma}\geq\frac{1}{3}$ for $\gamma\geq 6$,
$B_{n}(\hat{\theta}_{\tau,p})=O_{p}\left(\frac{\sqrt{\log n}}{n^{1/4}}\right)\quad\text{uniformly in }\tau\in[\varepsilon,1-\varepsilon].$
Step 2. Consider the estimator
$\hat{\xi}_{\tau}^{(2)}$ $\triangleq\hat{\xi}_{\tau}-\frac{G^{-1}(\theta_{\tau})B_{n}(\hat{\theta}_{\tau,p})}{\sqrt{n}},$
where, by Step 1, ${G^{-1}(\theta_{\tau})B_{n}(\hat{\theta}_{\tau,p})}/{\sqrt{n}}=O_{p}\left({n^{-3/4}\sqrt{\log n}}\right)$.
By Lemma A.2, we get
$\hat{g}_{\tau}(\hat{\xi}_{\tau}^{(2)})$ $=\frac{1}{\sqrt{n}}B^{\circ}_{n}(\theta_{\tau})+(\tau\mathbb{E}Z-\tau\mathbb{E}_{n}Z)+\frac{1}{\sqrt{n}}B_{n}(\hat{\xi}_{\tau}^{(2)})$
$+G(\theta_{\tau})(\hat{\xi}_{\tau}^{(2)}-\theta_{\tau})+(\hat{\xi}_{\tau}^{(2)}-\theta_{\tau})^{\prime}\frac{\partial G(\theta_{\tau})}{\partial\theta}(\hat{\xi}_{\tau}^{(2)}-\theta_{\tau})+O_{p}\left(\frac{1}{n^{3/2}}\right).$ (19)
Then, by definition of $\hat{\xi}_{\tau}^{(2)}$,
$\hat{g}_{\tau}(\hat{\xi}_{\tau}^{(2)})=$ $\frac{B_{n}(\hat{\xi}_{\tau}^{(2)})-B_{n}(\hat{\theta}_{\tau,p})}{\sqrt{n}}+(\hat{\xi}_{\tau}^{(2)}-\theta_{\tau})^{\prime}\frac{\partial G(\theta_{\tau})}{\partial\theta}(\hat{\xi}_{\tau}^{(2)}-\theta_{\tau})+O_{p}\left(\frac{1}{n^{3/2}}\right).$ (20)
Define $r_{n}$ to be a sequence satisfying $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau,p}-\hat{\xi}_{\tau}^{(2)}\|=O_{p}(r_{n}^{-1})$ (Lemma A.6 implies uniform consistency of $\hat{\theta}_{\tau,p}$ and that $r_{n}$ can be taken to be at least $\sqrt{n}$).
By Lemma A.3,
$B_{n}(\hat{\xi}_{\tau}^{(2)})-B_{n}(\hat{\theta}_{\tau,p})=O_{p}\left(\sqrt{\frac{{\log r_{n}}}{r_{n}}}\right)+o_{p}\left(\frac{\log r_{n}}{n^{\frac{\gamma-2}{2\gamma}}}\right),$
Then (20) becomes
$\hat{g}_{\tau}(\hat{\xi}_{\tau}^{(2)})=O_{p}\left(\frac{\sqrt{\log n}}{\sqrt{nr_{n}}}\right)+o_{p}\left(\frac{\log n}{{n^{1-\frac{1}{\gamma}}}}\right),$ (21)
where we replaced $\log r_{n}$ with the faster growing sequence $\log\sqrt{n}=O(\log n)$.
By Lemma A.2 applied to $\hat{\theta}_{\tau,p}$ and the definition of $\hat{\xi}_{\tau}$,
$\hat{\theta}_{\tau,p}$ $=\hat{\xi}_{\tau}+G^{-1}(\theta_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau,p})-\frac{G^{-1}(\theta_{\tau})B_{n}(\hat{\theta}_{\tau,p})}{\sqrt{n}}$
$-G^{-1}(\theta_{\tau})(\hat{\theta}_{\tau,p}-\theta_{\tau})^{\prime}\frac{\partial G(\theta_{\tau})}{\partial\theta}(\hat{\theta}_{\tau,p}-\theta_{\tau})+O_{p}\left(\frac{1}{n^{3/2}}\right).$
So by (19) and the definition of $\hat{\xi}_{\tau}^{(2)}$, we get
$\hat{\theta}_{\tau,p}-\hat{\xi}_{\tau}^{(2)}=G^{-1}(\theta_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau,p})+O_{p}(n^{-1}),$
which implies we can take $r_{n}^{-1}$ as the rate of $\hat{g}_{\tau}(\hat{\theta}_{\tau,p})=O_{p}(r_{n}^{-1})$.
Step 3. By definition of $\hat{\theta}_{\tau,p}$, $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}({\hat{\theta}_{\tau,p}})\|_{p}\leq\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}({\hat{\xi}_{\tau}^{(2)}})\|_{p}$. Then from (21), we obtain
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}({\hat{\theta}_{\tau,p}})\|_{p}\leq\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}({\hat{\xi}_{\tau}^{(2)}})\|_{p}=O_{p}\left(\frac{\sqrt{\log n}}{\sqrt{nr_{n}}}\right)+o_{p}\left(\frac{\log n}{{n^{1-\frac{1}{\gamma}}}}\right).$
On the right-hand side of this inequality, suppose that the first term dominates the second term. Then we have
$(r_{n}^{-1})^{\frac{1}{2}}$ $=O\left(\frac{\sqrt{\log n}}{n^{\frac{1}{2}-\frac{1}{\gamma}}}\right),$
or, equivalently,
$r_{n}^{-1}$ $=O\left(\frac{\log n}{n^{1-\frac{2}{\gamma}}}\right).$
By Step 2, it implies the statement of the lemma.
Suppose, instead, that the second term dominates the first term. Then by Step 2, $r_{n}^{-1}=O\left(n^{-1+\frac{1}{\gamma}}\log n\right)$. The statement of the lemma follows.
A.3 Proof of Theorem 1
The proof of Theorem 1 summarizes the results of auxiliary lemmas in Appendices A.1 and A.2.
Step 1 (Uniform consistency). The estimators under consideration are uniformly consistent, $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|=O_{p}\left({1}/{\sqrt{n}}\right)$. Specifically, the QR estimator is analyzed in Lemma A.5; the exact IVQR estimator is analyzed in Lemma A.6.
Step 2 (Generic stochastic expansion). By Step 1, we can apply Lemma A.2 with $r_{n}=\sqrt{n}$ to obtain
$\hat{\theta}_{\tau}-G^{-1}\hat{g}_{\tau}(\hat{\theta}_{\tau})=\hat{\xi}_{\tau}-G^{-1}\left[\frac{B_{n}(\hat{\theta}_{\tau})}{\sqrt{n}}+\frac{1}{2}(\hat{\theta}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau}-\theta_{\tau})\right]+O_{p}\left(\frac{1}{n^{3/2}}\right).$
Step 3 (Bounds on remainder in asymptotic linear expansion). Now we can use the remaining lemmas to bound the orders of the terms in the expansion. The result in equation (6) follows from Lemma A.3 with $\hat{\theta}_{\tau}^{\ast}=\theta_{\tau}$ and $r_{n}=\sqrt{n}$. The first equation in (5) is stated in Lemma A.5. Similarly, Lemma A.7 yields the second equation in (5). As a result, we have
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\hat{\xi}_{\tau}\|=O_{p}\left(\frac{\sqrt{\log n}}{n^{3/4}}\right),$ (22)
which is a uniform Bahadur-Kiefer expansion for both QR and exact IVQR estimators.
Step 4 (Analysis of quadratic term). The empirical process $\sqrt{n}(\hat{\xi}_{\tau}-\theta_{\tau})$ indexed by $\tau\in[\varepsilon,1-\varepsilon]$ is Donsker (see Step 2 of Lemma A.6), which implies (7), i.e.,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\xi}_{\tau}-\theta_{\tau}\|=O_{p}\left(\frac{1}{n^{1/2}}\right).$ (23)
Then by Step 3, we have, uniformly in $\tau\in[\varepsilon,1-\varepsilon]$,
$(\hat{\theta}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\theta}_{\tau}-\theta_{\tau})=(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})+O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right).$
Hence, the expansion in Step 2 becomes
$\hat{\theta}_{\tau}=\hat{\xi}_{\tau}+G^{-1}\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\frac{B_{n}(\hat{\theta}_{\tau})}{\sqrt{n}}-\frac{1}{2}(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})\right]+R_{n,\tau},$
with $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|R_{n,\tau}\|=O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right)$.
Appendix B Second-order bias correction, proofs
B.1 Auxiliary results
Lemma B.1.
Consider any random sequence $\hat{\theta}\in\Theta$ a.s. Under Assumption 3, the following expectations exist,
$\mathbb{E}\|\hat{g}_{\tau}(\hat{\theta})\|=O(1),$ (24)
$\mathbb{E}\|B^{\circ}_{n}(\hat{\theta})\|=O(1).$ (25)
Proof.
By the triangular inequality,
$\|\hat{g}_{\tau}(\hat{\theta})\|\leq\frac{1}{n}\sum_{i=1}^{n}\|Z_{i}\|(1+\tau)$ (26)
Therefore, $\mathbb{E}\|\hat{g}_{\tau}(\hat{\theta})\|\leq 2\mathbb{E}\|Z\|.$ Since the right-hand side does not depend on $n$, equation (24) holds.
By definition, we have
$B^{\circ}_{n}(\theta)\triangleq\sqrt{n}(\mathbb{E}_{n}1\{Y\leq W^{\prime}\theta\}Z-\mathbb{E}1\{Y\leq W^{\prime}\theta\}Z).$ (27)
We can bound $\mathbb{E}\|B^{\circ}_{n}(\hat{\theta})\|$ using a maximal inequality for an appropriately chosen empirical process. Consider the functions
$f_{\theta,\alpha}:\,(y,w,z)\mapsto 1\{y-w^{\prime}\theta\leq 0\}\alpha^{\prime}z,$
and the corresponding function class
$\mathcal{F}$ =$=\bigg{\{}f_{\theta,\alpha}:\,\,\theta\in\Theta,\|\alpha\|=1\bigg{\}}.$
Note that $\mathbb{G}_{n}f_{\theta,e_{j}}=B^{\circ}_{n}(\hat{\theta})^{\prime}e_{j}$, where $e_{j}\triangleq(0,\dots 0,1,0,\dots 0)^{\prime}$ with 1 in $j$-th position. By Assumption 3, $\mathcal{F}$ admits an envelope $F(y,w,z)\equiv\|z\|$.
We use the maximal inequality in Ota et al. (2019) to establish (25). To do so, we verify the three conditions of this lemma. First,
$\mathbb{E}f_{\theta,\alpha}^{2}(Y,W,Z)$ $=\mathbb{E}\left[1\big{\{}Y-W^{\prime}\theta\leq 0\big{\}}\alpha^{\prime}Z\right]^{2}\leq\mathbb{E}\|Z\|^{2}.$
Therefore, the variance parameter of the process is
$\sigma_{n}^{2}\triangleq\sup_{f\in\mathcal{F}}\mathbb{E}f^{2}(Y,W,Z)\leq\mathbb{E}\|Z\|^{2}.$
Second, using Lemma 2 in Ota et al. (2019) we have
$\mathbb{E}\max_{1\leq i\leq n}F^{2}(Y_{i},W_{i},Z_{i})=\mathbb{E}\max_{1\leq i\leq n}\|Z_{i}\|^{2}=o(n^{2/\gamma}).$
Third, because the function class $\mathcal{F}$ is a VC class with the envelope $\|Z\|$, there exist constants $A$ and $V$ independent of $n$ such that the standard entropy bound
$\sup_{Q}N\left(\mathcal{F},\|\cdot\|_{Q,2},\eta\|Z\|_{Q,2}\right)\leq(A/\eta)^{V}\text{ for all }\eta\in(0,1]$
holds (e.g., van der Vaart and Wellner, 1996, Section 2.6). Here the supremum is taken with respect to all finitely discrete measures $Q$ and $\|\cdot\|_{Q,2}$ is the $L^{2}(Q)$ norm.
Finally, applying Lemma 1 in Ota et al. (2019), we obtain
$\mathbb{E}\sup_{\theta\in\Theta,\|\alpha\|=1}\|\mathbb{G}_{n}f_{\theta,\alpha}\|\lesssim\sqrt{V\sigma_{n}^{2}\log(Am/\sigma_{n})}+\frac{V\sqrt{\mathbb{E}\max_{1\leq i\leq n}\|Z_{i}\|^{2}}}{\sqrt{n}}\log(Am/\sigma_{n})$
$=O\left(1\right)+o\left(\frac{1}{n^{\frac{\gamma-2}{2\gamma}}}\right),$ (28)
It follows that
$\|B^{\circ}_{n}(\hat{\theta})\|\leq\sup_{\theta\in\Theta,\|\alpha\|=1}\|\mathbb{G}_{n}f_{\theta,\alpha}\|$
which implies
$\mathbb{E}\|B^{\circ}_{n}(\hat{\theta})\|\leq\mathbb{E}\sup_{\theta\in\Theta,\|\alpha\|=1}\|\mathbb{G}_{n}f_{\theta,\alpha}\|=O\left(1\right).$
Lemma B.2.
Consider $\hat{\theta}_{n}$ such that $\hat{\theta}_{n}\in\Theta$ and $\hat{\theta}_{n}-\theta_{\tau}=o_{p}(1)$. Then $\mathbb{E}\|\hat{\theta}_{n}-\theta_{\tau}\|^{q}=o(1)$ for any $q>0$.
Proof.
Notice that
$\|\hat{\theta}_{n}-\theta_{\tau}\|^{q}\leq(\max_{\theta\in\Theta}\|\theta-\theta_{\tau}\|)^{q}\leq\text{diam}(\Theta)^{q}.$
Hence the sequence $\|\hat{\theta}_{n}-\theta_{\tau}\|^{q}$ is uniformly integrable. By Proposition 4.12 from Kallenberg (2006), $\mathbb{E}\|\hat{\theta}_{n}-\theta_{\tau}\|^{q}\ =o(1)$.
Lemma B.3.
Consider $\hat{\theta}=\hat{\theta}_{\tau,p}$ defined in (3) for some $p\in[1,\infty]$ or $\hat{\theta}=\hat{\theta}_{\tau,\text{QR}}$, where $\tau\in(0,1)$. Under Assumptions 1, 2, and 3, $B_{n}(\hat{\theta})=B^{1}_{n}+B^{2}_{n}$, where the two components satisfy
$\mathbb{E}\frac{1}{\sqrt{n}}B^{1}_{n}(\hat{\theta})=\mathbb{E}\left(\frac{\hat{g}_{\tau}(\hat{\theta})+\hat{g}_{\tau}^{*}(-\hat{\theta})}{2}\right)+\frac{1}{n}\kappa(\tau),$
with
$\kappa(\tau)\triangleq\mathbb{E}\left(\tau-\frac{1}{2}\right)f_{\varepsilon}(0|W,Z)ZW^{\prime}G^{-1}Z,$
while $B^{2}_{n}=O_{p}\left(n^{-3/4}\sqrt{\log n}\right)$ and $n^{-1/2}B^{2}_{n}$ is uniformly integrable.
Proof.
The proof proceeds in six steps.
Step 1. By Lemma B.1, $\mathbb{E}B_{n}(\hat{\theta})$ exists. Note that
$\frac{1}{\sqrt{n}}\mathbb{E}B_{n}(\hat{\theta})$ $=\frac{1}{\sqrt{n}}\mathbb{E}\left(B^{\circ}_{n}(\hat{\theta})-B^{\circ}_{n}(\theta_{\tau})\right)$
$=\mathbb{E}\left(1\{Y\leq W^{\prime}\hat{\theta}\}Z\right)-\mathbb{E}g^{\circ}(\hat{\theta}).$ (29)
Theorem 1 implies
$\hat{\theta}=\theta_{\tau}-\frac{1}{n}G^{-1}\sum_{i=1}^{n}\left(1\{Y_{i}\leq W_{i}^{\prime}\theta_{\tau}\}-\tau\right)Z_{i}+\tilde{R}_{n},$ (30)
where $\tilde{R}_{n}=O_{p}\left(n^{-3/4}\sqrt{\log{n}}\right)$. Since by construction, $\hat{\theta}$ is restricted to a compact set $\Theta$, it is bounded. The term $\left(1\{Y_{i}\leq W_{i}^{\prime}\theta_{\tau}\}-\tau\right)Z_{i}$ has bounded moments up to order $\gamma$ by Assumption 3. As a result, the remainder term $\tilde{R}_{n}$ has bounded moments up to order $\gamma$.
Step 2. Define $\hat{\varepsilon}_{i}\triangleq Y_{i}-W_{i}^{\prime}\hat{\theta}$ and split the first term in equation (29) as follows:
$\mathbb{E}1\{Y_{i}\leq W_{i}^{\prime}\hat{\theta}\}Z_{i}=\mathbb{E}1\{\hat{\varepsilon}_{i}=0\}Z_{i}+\mathbb{E}1\{\hat{\varepsilon}_{i}<0\}Z_{i}.$ (31)
We can use (30) to isolate an influence of observation $i$, $\lambda_{i}\triangleq-W_{i}^{\prime}G^{-1}Z_{i}(1\{Y_{i}\leq W_{i}^{\prime}\theta_{\tau}\}-\tau)$. Without loss of generality for i.i.d. data, we consider $i=1$. The indicator $1\{\hat{\varepsilon}_{1}<0\}$ can be rewritten as $1\left\{Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}+n^{-1}\lambda_{1}\right\},$ where
$\hat{\theta}_{-1}\triangleq\hat{\theta}+\frac{1}{n}G^{-1}\left(1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\}-\tau\right)Z_{1}=\theta_{\tau}-\frac{1}{n}G^{-1}\sum_{j=2}^{n}\left(1\{Y_{j}\leq W_{j}^{\prime}\theta_{\tau}\}-\tau\right)Z_{j}+\tilde{R}_{n}$
is equal to $\hat{\theta}$ without the linear influence of the observation $i=1$.
Then, using Taylor’s theorem (justified below equation (32)),
$\mathbb{E}\left[Z_{1}P\left(Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}+\frac{1}{n}\lambda_{1}|1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\},Z_{1},W_{1}\right)\right]$
$=\mathbb{E}\left[Z_{1}P\left(Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\},Z_{1},W_{1}\right)\right]$
$+\mathbb{E}\left[\frac{1}{n}Z_{1}\lambda_{1}f_{Y_{1}}(W_{1}^{\prime}\hat{\theta}_{-1}|1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\},Z_{1},W_{1})\right]+\frac{1}{n^{2}}\psi^{1}_{n},$
where by the mean value theorem $\psi^{1}_{n}=O(1)$.
Note also that the first term in the Taylor expansion can be rewritten as
$\mathbb{E}\left[Z_{1}P(Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\},Z_{1},W_{1})\right]=\mathbb{E}\left[Z_{1}1\{Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}\}\right]=\mathbb{E}\left[Z_{1}P(Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|Z_{1},W_{1})\right].$
The following argument justifies the use of Taylor’s theorem here. The function $f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},W_{1},Z_{1})$ is measurable as a limit of measurable functions (increments of conditional CDF). Therefore, for any non-negative measurable function $\phi(W_{1},Z_{1})$ with finite expectation, the integral
$\mathbb{E}\left[\phi(W_{1},Z_{1})f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},W_{1},Z_{1})\right]$ exists (but may take infinite values). By the law of iterated expectations,
$\mathbb{E}\left[\phi(W_{1},Z_{1})f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},W_{1},Z_{1})\right]=\mathbb{E}\left[\phi(W_{1},Z_{1})f_{Y_{1}}(y|W_{1},Z_{1})\right]$
(see Step 5 below for a detailed justification based on Fubini-Tonelli theorem). By Assumption 2, $f_{Y_{1}}(y|W_{1},Z_{1})$ is uniformly bounded and
$\mathbb{E}\left[\phi(W_{1},Z_{1})f_{Y_{1}}(y|W_{1},Z_{1})\right]\leq\bar{f}\cdot\mathbb{E}\left[\phi(W_{1},Z_{1})\right]<\infty.$ (32)
The same is true for the derivative of the density $\partial f_{Y_{1}}$ in place of $f_{Y_{1}}$, by Assumption 2. Therefore, $P(f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},W_{1},Z_{1})=\infty)=0$ and $P(\partial f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},W_{1},Z_{1})=\infty)=0$, which justifies the Taylor expansion of the expectations of the conditional PDF above. By this property (a.s. smoothness of $f_{Y_{1}}(y|\hat{\theta}_{-1},\lambda_{1},Z_{1},W_{1})$) and equation (30),
$\mathbb{E}\left[Z_{1}\lambda_{1}f_{Y_{1}}(W_{1}^{\prime}\hat{\theta}_{-1}|\hat{\theta}_{-1},W_{1},\lambda_{1},Z_{1})\right]$ $=\mathbb{E}\left[Z_{1}\lambda_{1}f_{Y_{1}}(W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1},\lambda_{1})\right]$
$+\mathbb{E}\left(Z_{1}\lambda_{1}W_{1}^{\prime}(\hat{\theta}_{-1}-\theta_{\tau})\partial f_{Y_{1}}(\xi|W_{1},Z_{1},\lambda_{1})\right),$
where $\xi$ is some random variable that takes values between $W_{1}^{\prime}\hat{\theta}_{-1}$ and $W_{1}^{\prime}\theta_{\tau}$. By the boundedness of $\hat{\theta}_{-1}\in\Theta$, Assumption 3, the bound on the derivative of the density in Assumption 2, and the fact that $\hat{\theta}_{-1}=\theta_{\tau}+O_{p}\left({1}/{\sqrt{n}}\right)$, these expectations exist and the second term denoted as
$\psi^{2}_{n}\triangleq Z_{1}\lambda_{1}W_{1}^{\prime}(\hat{\theta}_{-1}-\theta_{\tau})\partial f_{Y_{1}}(\xi|W_{1},Z_{1},\lambda_{1})$
is of order $O_{p}\left({1}/{\sqrt{n}}\right)$.
By the definition of $\lambda_{1}$, the first term can be rewritten as
$\mathbb{E}\left[Z_{1}\lambda_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\lambda_{1})\right]=$ $-\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}1\{\varepsilon_{i}\leq 0\}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\lambda_{1})\right]$
$+\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}\tau f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right].$
Finally, (31) becomes
$\mathbb{E}\left(1\{Y_{i}\leq W_{i}^{\prime}\hat{\theta}\}Z_{i}-\frac{1}{n}\psi^{2}_{n}\right)=\mathbb{E}\left[Z_{1}P(Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|Z_{1},W_{1})\right]$
$+\frac{\tau}{n}\mathbb{E}\left[f_{\varepsilon_{1}}(0|W_{1},Z_{1})Z_{1}W_{1}^{\prime}G^{-1}Z_{1}\right]+\mathbb{E}\left[1\{\hat{\varepsilon}_{1}=0\}Z_{1}\right]+\frac{1}{n}\Xi_{\tau}+\psi_{n}^{1},$ (33)
where the term $\Xi_{\tau}\triangleq-\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}1\{\varepsilon_{1}\leq 0\}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\lambda_{1})\right]$ and $\frac{1}{n}\psi^{2}_{n}=O_{p}\left(n^{-3/2}\right).$
Step 3. Now consider $\mathbb{E}g^{\circ}(\hat{\theta})$, the second term in (29). Let $(Y_{n+1},W_{n+1},Z_{n+1})$ be a copy of $(Y,W,Z)$, which is independent of the sample $\{Y_{i},W_{i},Z_{i}\}_{i=1}^{n}$. Also, define
$\lambda_{n+1,1}\triangleq-\frac{1}{n}W_{n+1}^{\prime}G^{-1}Z_{1}(1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\}-\tau),$
which satisfies $\mathbb{E}\lambda_{n+1,1}=0$. Then
$\mathbb{E}\left(g^{\circ}(\hat{\theta})-\frac{1}{n}\psi^{4}_{n}\right)$ $=\mathbb{E}(1\{Y_{n+1}\leq W_{n+1}^{\prime}\hat{\theta}\}Z_{n+1}-\frac{1}{n}\psi^{4}_{n})$
$=\mathbb{E}\left(P\{Y_{n+1}\leq W_{n+1}^{\prime}\hat{\theta}_{-1}-\frac{1}{n}\lambda_{n+1,1}|W_{n+1},Z_{n+1}\}Z_{n+1}-\frac{1}{n}\psi^{4}_{n}\right)$
$=\mathbb{E}\left(P\{Y_{n+1}<W_{n+1}^{\prime}\hat{\theta}_{-1}|W_{n+1},Z_{n+1}\}Z_{n+1}\right)$
$+\frac{1}{n}\mathbb{E}\left(Z_{n+1}\lambda_{n+1,1}f_{Y_{n+1}}(W_{n+1}^{\prime}\theta_{\tau}|1\{Y_{1}\leq W_{1}^{\prime}\theta_{\tau}\},Z_{n+1},W_{n+1})\right)$ (34)
$+\frac{1}{n^{2}}\psi_{n}^{3},$
where $\psi_{n}^{3}=O(1)$ by the mean value theorem and
$\psi^{4}_{n}\triangleq Z_{n+1}\lambda_{n+1}W_{n+1}^{\prime}(\hat{\theta}_{-1}-\theta_{\tau})\partial f_{Y_{n+1}}(\xi|W_{n+1},Z_{n+1},\lambda_{n+1}).$
The rate $\psi^{4}_{n}=O_{p}(n^{-1/2})$ is derived by an argument similar to the one below equation (32). Note that the term in line (34) is equal to zero since $\mathbb{E}\left(\lambda_{n+1,1}|Y_{n+1},W_{n+1},Z_{n+1}\right)=0$ by the i.i.d. data assumption. Combining this equality with (33) yields
$\mathbb{E}\left(1\{Y_{1}\leq W_{1}^{\prime}\hat{\theta}\}Z_{1}\right)-\mathbb{E}g^{\circ}(\hat{\theta})-\frac{1}{n}\mathbb{E}(\psi^{2}_{n}-\psi^{4}_{n})-\frac{\psi_{n}^{1}-\psi_{n}^{3}}{n^{2}}$
$=\mathbb{E}\left[Z_{1}P\{Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|W_{1},Z_{1}\}\right]-\mathbb{E}\left[Z_{n+1}P\{Y_{n+1}<W_{n+1}^{\prime}\hat{\theta}_{-1}|W_{n+1},Z_{n+1}\}\right]$
$+\frac{\tau}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right]+\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}+\frac{1}{n}\Xi_{\tau}.$ (35)
Step 4. Let us simplify the first two terms of equation (35). Define
$\hat{\zeta}_{-1}\triangleq-\frac{1}{n}G^{-1}\sum_{j=2}^{n}\left(1\{\varepsilon_{j}\leq 0\}-\tau\right)Z_{j},$
so that $\hat{\zeta}_{-1}$ has zero mean and is independent of $Y_{1}$ and $\hat{\theta}_{-1}=\theta_{\tau}+\hat{\zeta}_{-1}+\tilde{R}_{n}$.
Denote $\hat{\xi}_{1}\triangleq Y_{1}-W_{1}^{\prime}\hat{\zeta}_{-1}$. Apply Taylor’s theorem (as in Step 2) to obtain
$\mathbb{E}\left[Z_{1}P\{Y_{1}<W_{1}^{\prime}\hat{\theta}_{-1}|W_{1},Z_{1},W_{1}^{\prime}\tilde{R}_{n}\}\right]$
$=\mathbb{E}\left[Z_{1}P\{\hat{\xi}_{1}<W_{1}^{\prime}(\theta_{\tau}+\tilde{R}_{n})|W_{1},Z_{1},W_{1}^{\prime}\tilde{R}_{n}\}\right]$
$=\mathbb{E}\left[Z_{1}P(\hat{\xi}_{1}<W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1})\right]+\mathbb{E}\left[Z_{1}W_{1}^{\prime}\tilde{R}_{n}f_{\hat{\xi}_{1}}(W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1},\tilde{R}_{n})\right]$
$+\frac{1}{2}\mathbb{E}\left[Z_{1}\tilde{R}_{n}^{\prime}W_{1}\partial f_{\hat{\xi}_{1}}(\eta|W_{1},Z_{1},\tilde{R}_{n})W_{1}^{\prime}\tilde{R}_{n}\right],$ (36)
where $\eta$ is a random scalar that takes values between $W_{1}^{\prime}\theta_{\tau}$ and $W_{1}^{\prime}\tilde{R}_{n}$. By Step 1, $\tilde{R}_{n}=O_{p}\left(n^{-3/4}\sqrt{\log n}\right)$ has a finite second moment. Therefore, the last term in (36) is finite.
For the second term in (36), note that
$\tilde{R}_{n}f_{\hat{\xi}_{1}}(W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1},\tilde{R}_{n})$ $=\tilde{R}_{n}f_{\varepsilon_{1}}(W_{1}^{\prime}\hat{\zeta}_{-1}|W_{1},Z_{1},\tilde{R}_{n})$
$=\tilde{R}_{n}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\tilde{R}_{n})+\partial f_{\varepsilon_{1}}(\tilde{\eta}|W_{1},Z_{1},\tilde{R}_{n})\tilde{R}_{n}W_{1}^{\prime}\hat{\zeta}_{-1}$
$=\tilde{R}_{n}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\tilde{R}_{n})+O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right),$
where $\tilde{\eta}$ is a random scalar that takes values between 00 and $W_{1}^{\prime}\hat{\zeta}_{-1}$. The last equality follows since $\tilde{R}_{n}=O_{p}\left(n^{-3/4}\sqrt{\log n}\right)$ by Step 1 and $\hat{\zeta}_{-1}=O_{p}\left(1/{\sqrt{n}}\right)$. As before, let us introduce
$\psi^{5}_{n}\triangleq\partial f_{\varepsilon_{1}}(\tilde{\eta}|W_{1},Z_{1},\tilde{R}_{n})\tilde{R}_{n}W_{1}^{\prime}\hat{\zeta}_{-1}+\frac{1}{2}Z_{1}\tilde{R}_{n}^{\prime}W_{1}\partial f_{\hat{\xi}_{1}}(\tilde{\eta}|W_{1},Z_{1},\tilde{R}_{n})W_{1}^{\prime}\tilde{R}_{n}=O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right).$
By the boundedness of $\gamma$-moments of $\tilde{R}_{n}$ (Assumption 3), $\mathbb{E}\psi^{5}_{n}$ exists. Hence, (36) becomes
$\mathbb{E}\left[Z_{1}P(Y_{1}-W_{1}^{\prime}\hat{\zeta}_{-1}<W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1})\right]+\mathbb{E}\left[Z_{1}W_{1}^{\prime}\tilde{R}_{n}f_{\varepsilon_{1}}(0|W_{1},Z_{1},\tilde{R}_{n})\right]+\mathbb{E}\psi^{5}_{n}.$
Similarly, using the i.i.d. assumption,
$\mathbb{E}\left[Z_{n+1}P\{Y_{n+1}<W_{n+1}^{\prime}\hat{\theta}_{-1}|W_{n+1},Z_{n+1}\}\right]$
$=\mathbb{E}\left[Z_{n+1}P(Y_{n+1}-W_{n+1}^{\prime}\hat{\zeta}_{-1}<W_{n+1}^{\prime}\theta_{\tau}|W_{n+1},Z_{n+1})\right]$
$+\mathbb{E}\left[Z_{n+1}W_{n+1}^{\prime}\tilde{R}_{n}f_{\varepsilon_{n+1}}(W_{n+1}^{\prime}\hat{\zeta}_{-1}|W_{n+1},Z_{n+1},\tilde{R}_{n})\right]+\mathbb{E}\psi^{6}_{n}$
$=\mathbb{E}\left[Z_{1}P(Y_{1}-W_{1}^{\prime}\hat{\zeta}_{-1,n}<W_{1}^{\prime}\theta_{\tau}|W_{1},Z_{1})\right]+\mathbb{E}\left[f_{\varepsilon_{1}}(0|W_{1},Z_{1})Z_{1}W_{1}^{\prime}\right]\cdot\mathbb{E}\tilde{R}_{n}+\mathbb{E}\psi^{6}_{n},$
where
$\psi^{6}_{n}\triangleq\partial f_{\varepsilon_{1}}(\tilde{\eta}|W_{1},Z_{1})\tilde{R}_{n}W_{1}^{\prime}\hat{\zeta}_{-1}+\frac{1}{2}Z_{1}\tilde{R}_{n}^{\prime}W_{1}\partial f_{\hat{\xi}_{1}}(\tilde{\eta}|W_{1},Z_{1})W_{1}^{\prime}\tilde{R}_{n}=O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right).$
To summarize, (35) becomes
$\mathbb{E}\left(1\{Y_{1}\leq W_{1}^{\prime}\hat{\theta}\}Z_{1}\right)-\mathbb{E}g^{\circ}(\hat{\theta})-\mathbb{E}\left(\frac{\psi^{1}_{n}-\psi^{3}_{n}}{n^{2}}+\frac{\psi_{n}^{2}-\psi^{4}_{n}}{n}+\psi_{n}^{5}-\psi^{6}_{n}\right)$
$=\frac{\tau}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right]+\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}+\frac{1}{n}\Xi_{\tau}$
$+\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})(f_{\varepsilon_{1}}(0|W_{1},Z_{1},\tilde{R}_{n})-f_{\varepsilon_{1}}(0|W_{1},Z_{1}))\right].$ (37)
Step 5. Let us study the last term in equation (37). For any $t\geq 0$, consider an auxiliary function
$\Psi(t)$ $\triangleq\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})(F_{\varepsilon_{1}}(t|W_{1},Z_{1},\tilde{R}_{n})-F_{\varepsilon_{1}}(t|W_{1},Z_{1}))\right].$
By definition, for every $t\geq 0$,
$\Psi(t)=\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})1\{0<\varepsilon_{1}\leq t\}\right]-\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})1\{0<\varepsilon_{1}\leq t\}\right]=0.$
By the existence of the corresponding conditional PDF (possibly taking infinite values),
[…]
By the Fubini-Tonelli theorem for product measures, we can exchange the order of integration,
[…]
Hence, by the main theorem of calculus, for all $e\geq 0$,
$\frac{\partial\Psi(e)}{\partial e}=\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})(f_{\varepsilon_{1}}(e|W_{1},Z_{1},\tilde{R}_{n})-f_{\varepsilon_{1}}(e|W_{1},Z_{1}))\right].$
Since the function $\Psi(t)\equiv 0$, we have
$\frac{\partial\Psi(0)}{\partial e}=\mathbb{E}\left[Z_{1}W_{1}^{\prime}(\tilde{R}_{n}-\mathbb{E}\tilde{R}_{n})(f_{\varepsilon_{1}}(0|W_{1},Z_{1},R_{n})-f_{\varepsilon_{1}}(0|W_{1},Z_{1}))\right]=0.$
Therefore, equation (37) becomes
$\mathbb{E}\left(1\{Y_{1}\leq W_{1}^{\prime}\hat{\theta}\}Z_{1}\right)-\mathbb{E}g^{\circ}(\hat{\theta})-\mathbb{E}\psi_{n}$
$=\frac{\tau}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right]+\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}+\frac{1}{n}\Xi_{\tau},$ (38)
where
$\psi_{n}\triangleq\left(\frac{\psi^{1}_{n}-\psi^{3}_{n}}{n^{2}}+\frac{\psi_{n}^{2}-\psi^{4}_{n}}{n}+\psi_{n}^{5}-\psi^{6}_{n}\right)=O_{p}\left(\frac{\sqrt{\log n}}{n^{5/4}}\right).$
Step 6. Let us simplify the second and the third terms in equation (38). The latter can be rewritten as
$\mathbb{E}\left(1\{Y_{1}\leq W_{i}^{\prime}\hat{\theta}\}Z_{1}\right)-\mathbb{E}\left(1\{Y_{n+1}\leq W_{n+1}^{\prime}\hat{\theta}\}Z_{n+1}\right)$ (39)
$=\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}-\mathbb{E}\left(1\{Y_{1}\geq W_{i}^{\prime}\hat{\theta}\}Z_{1}\right)+\mathbb{E}\left(1\{Y_{n+1}\geq W_{n+1}^{\prime}\hat{\theta}\}Z_{n+1}\right)$
$=\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}-\mathbb{E}\left(1\{-Y_{1}\leq W_{i}^{\prime}(-\hat{\theta})\}-(1-\tau)\right)Z_{1}+\mathbb{E}\left(1\{-Y_{n+1}\leq W_{n+1}^{\prime}(-\hat{\theta})\}-(1-\tau)\right)Z_{n+1}$ (40)
$=\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}-\bigg{[}\frac{1}{n}\mathbb{E}\left[Z_{1}W_{i}^{\prime}G^{-1}Z_{1}(1-\tau)f_{\varepsilon}(0|W_{1},Z_{1})\right]+\frac{1}{n}\Xi^{*}_{\tau}+\mathbb{E}1\{-\hat{\varepsilon}_{1}=0\}Z_{1}\bigg{]}+\mathbb{E}\psi^{*}_{n}$
$=\bigg{[}\frac{1}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}(\tau-1)f_{\varepsilon}(0|W_{1},Z_{1})\right]-\frac{1}{n}\Xi^{*}_{\tau}\bigg{]}+\mathbb{E}\psi^{*}_{n},$ (41)
where $\Xi^{*}_{\tau}\triangleq-\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}1\{-\varepsilon_{1}\leq 0\}f_{-\varepsilon_{1}}(0|W_{1},Z_{1},1\{-\varepsilon_{1}\leq 0\})\right]$, $\psi^{*}_{n}$ is the analog of $\psi_{n}$ corresponding to the moment condition for $-Y_{i}$, $-\theta$, and $(1-\tau)$, and the last equality uses (38). Notice that it follows from equations (39) and (40) that
$\mathbb{E}1\{\hat{\varepsilon}_{1}=0\}Z_{1}=\mathbb{E}\left(\hat{g}(\hat{\theta})+\hat{g}^{*}(-\hat{\theta})\right).$ (42)
Using the definition of $f_{\varepsilon_{1}}(0|W_{1},Z_{1},1\{\varepsilon_{1}\leq 0\})$ and Fubini-Tonelli theorem as in Step 5,
$\Xi_{\tau}$ $=-\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}1\{\varepsilon_{1}\leq 0\}f_{\varepsilon_{1}}(0|W_{1},Z_{1},1\{\varepsilon_{1}\leq 0\})\right]$
$=-\lim_{t\downarrow 0}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}1\{\varepsilon_{1}\leq 0\}\frac{1\{\varepsilon_{1}\leq 0\}-1\{\varepsilon_{1}\leq-t\}}{t}\right]$
$=-\lim_{t\downarrow 0}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}\frac{1\{\varepsilon_{1}\leq 0\}-1\{\varepsilon_{1}\leq-t\}}{t}\right]$
$=-\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right].$
Since $\varepsilon_{1}$ has conditional density by Assumption 2, the same argument can be applied to show that $\Xi^{*}_{\tau}=\Xi_{\tau}$. Hence, equations (38) and (41) imply
$\frac{1}{n}\Xi_{\tau}=-\frac{1}{2}\left(\mathbb{E}\left(\hat{g}(\hat{\theta})+\hat{g}^{*}(-\hat{\theta})\right)+\frac{1}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}f_{\varepsilon_{1}}(0|W_{1},Z_{1})\right]\right)+\frac{1}{2}(\mathbb{E}\psi_{n}-\mathbb{E}\psi^{*}_{n}).$
Finally, equations (38) and (42) yield
$\mathbb{E}\frac{1}{\sqrt{n}}(B_{n}(\hat{\theta})-B^{2}_{n})$ $=\frac{1}{n}\mathbb{E}\left[Z_{1}W_{1}^{\prime}G^{-1}Z_{1}\left(\tau-\frac{1}{2}\right)f_{\varepsilon}(0|W_{1},Z_{1})\right]+\mathbb{E}\left(\frac{\hat{g}(\hat{\theta})+\hat{g}^{*}(-\hat{\theta})}{2}\right),$
where $n^{-1/2}B^{2}_{n}\triangleq\frac{1}{2}(\psi_{n}+\psi^{*}_{n})$. By construction, $B^{2}_{n}=O_{p}\left(n^{-3/4}\sqrt{\log n}\right)$ and $n^{-1/2}B^{2}_{n}$ is uniformly integrable as the sum of uniformly integrable components. The proof is complete. ∎
Lemma B.4.
Suppose that a function $f(x)$ is four times continuously differentiable in a neighborhood of x. Then for sufficiently small $h\in\mathbb{R}$,
$\partial_{x}f(x)$ $=\frac{f(x+h)-f(x-h)}{2h}+O(h^{2}),$
$\partial_{x,x}f(x)$ $=\frac{f(x+h)-2f(x)+f(x-h)}{h^{2}}+O(h^{2}).$
Proof.
See Chapter 5 in Olver (2014) and p.884 in Abramowitz and Stegun (1972). ∎
B.2 Proofs of main results on bias correction
Proof of Theorem 2.
By Theorem 1,
$\hat{\theta}_{\tau}=\hat{\xi}_{\tau}+G^{-1}\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\frac{B_{n}(\hat{\theta}_{\tau})}{\sqrt{n}}-\frac{1}{2}(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})\right]+R_{n,\tau},$
where $R_{n,\tau}=O_{p}\left(n^{-5/4}\sqrt{\log n}\right)$, and $\|R_{n,\tau}\|$ is uniformly integrable by Lemmas B.1 and B.2. Lemma B.3 implies
$\frac{1}{\sqrt{n}}\mathbb{E}B^{1}_{n}(\hat{\theta}_{\tau})=\mathbb{E}\left(\frac{\hat{g}(\hat{\theta})+\hat{g}^{*}(-\hat{\theta})}{2}\right)+\frac{1}{n}\kappa(\tau).$
For correctly specified models, $\mathbb{E}\hat{\xi}_{\tau}=\theta_{\tau}$ and, for each component $j$, we have
$(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G_{j}(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})$ $=\mathbb{E}\hat{g}_{\tau}^{\prime}(\theta_{\tau})(G^{-1})^{\prime}\partial_{\theta}G_{j}(\theta_{\tau})G^{-1}\hat{g}_{\tau}(\theta_{\tau})$
$=\frac{1}{n}Q_{j}^{\prime}vec(\Omega).$
By definition of $\operatorname{Bias}(\hat{\theta}_{\tau})$, we can ignore the terms $R_{n,\tau}$ and $B_{n}^{2}$. The statement of the theorem follows. ∎
Proof of Theorem 3.
Theorem 1 implies the following asymptotic expansion for the bias-corrected estimator,
$\hat{\theta}_{bc}$ $=\hat{\xi}_{\tau}+G^{-1}\left[\hat{g}_{\tau}(\hat{\theta}_{\tau})-\frac{B_{n}(\hat{\theta}_{\tau})}{\sqrt{n}}-\frac{1}{2}(\hat{\xi}_{\tau}-\theta_{\tau})^{\prime}\partial_{\theta}G(\theta_{\tau})(\hat{\xi}_{\tau}-\theta_{\tau})\right]$ (43)
$-\mathbb{E}{G}^{-1}\frac{\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)}{2}+\frac{1}{n}G^{-1}\left[\kappa_{\tau}+\frac{1}{2}Q^{\prime}vec(\Omega)\right]$ (44)
$-(\hat{G}^{-1}-{G}^{-1})\frac{\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)}{2}+\frac{1}{n}(\hat{G}^{-1}-{G}^{-1})\left[\hat{\kappa}_{\tau}+\frac{1}{2}\hat{Q}^{\prime}vec(\hat{\Omega})\right]$ (45)
$+\frac{1}{n}G^{-1}\left[\hat{\kappa}_{\tau}-\kappa_{\tau}+\frac{1}{2}\hat{Q}^{\prime}vec(\hat{\Omega})-Q^{\prime}vec(\Omega)\right]+R_{n,\tau}$ (46)
$-{G}^{-1}\frac{\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)}{2}+\mathbb{E}{G}^{-1}\frac{\left(\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}^{*}(-\hat{\theta}_{\tau})\right)}{2}.$ (47)
By Theorem 2, the expectation of the sum of the terms in (43) and (44) is zero. By Theorem 1, Assumption 4, and the Mann-Wald and Delta theorems, the first term in (45) is
$o_{p}\left(\frac{1}{n^{1/3}\log n}\right)O_{p}\left(\frac{\log n}{n^{1-2/\gamma}}\right)=o_{p}\left(n^{-1}\right),$
since $\gamma\geq 6$. The same rate $o_{p}\left(n^{-1}\right)$ is true for the second term in (45) and the terms in (46) The last line, expression (47), has zero mean by Assumption 3. Therefore, $\operatorname{Bias}(\hat{\theta}_{bc})=0$. ∎
Proof of Lemma 1.
Notice that by the same arguments as in Lemma A.3, for any $h$,
$\mathbb{E}_{n}\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h\}}{2h}Z_{i}W_{j}-\mathbb{E}\frac{F_{Y}(W^{\prime}\theta+h|W,Z)}{2h}Z_{i}W_{j}\bigg{|}_{\theta=\hat{\theta}_{\tau}}$
$-\left(\mathbb{E}_{n}\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h\}}{2h}Z_{i}W_{j}-\mathbb{E}\frac{F_{Y}(W^{\prime}\theta-h|W,Z)}{2h}Z_{i}W_{j}\bigg{|}_{\theta=\hat{\theta}_{\tau}}\right)=O_{p}\left(\frac{\sqrt{\log h}}{\sqrt{nh}}\right)+o_{p}\left(\frac{\log h}{n^{5/6}h}\right).$
Then using Lemma B.4 for $h_{n}\to 0$ and the Delta theorem, we obtain
$\mathbb{E}_{n}\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{n}\}}{2h_{n}}Z_{i}W_{j}$
$=G_{i,j}(\theta)+O_{p}\left(\frac{1}{\sqrt{n}}\right)+O\left(h_{n}^{2}\right)+O_{p}\left(\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right)+o_{p}\left(\frac{\log h_{n}}{n^{5/6}h_{n}}\right).$
The overall rate is $\max\left\{n^{-1/2},h_{n}^{2},(nh_{n})^{-1/2}\sqrt{\log h_{n}},(n^{5/6}h_{n})^{-1}\log h_{n}\right\}$. By Lemmas A.3, A.5, A.7, and Assumption 2, this remainder rate is uniform in $\tau\in[\varepsilon,1-\varepsilon]$. Ignoring a logarithmic factor, we see that the bandwidth $h_{1,n}\propto{n^{-1/5}}$ delivers an optimal overall remainder rate $O_{p}\left(n^{-2/5}\sqrt{\log n}\right)$ that is uniform in $\tau\in[\varepsilon,1-\varepsilon]$
Similarly, by Lemma B.4,
$\widehat{({\partial_{\theta}G}_{i,\ell})}_{j}=e_{i}^{\prime}\partial_{\theta}G_{j}(\theta)e_{\ell}+O_{p}\left(\frac{1}{\sqrt{n}}\right)+O(h_{n}^{2})+O_{p}\left(\frac{\sqrt{\log h_{n}}}{\sqrt{n}\sqrt{h_{n}^{3}}}\right)+o_{p}\left(\frac{\log h_{n}}{n^{5/6}h_{n}^{2}}\right).$
Taking $h_{2,n}\propto n^{-1/7}$, we obtain the optimal remainder rate
$\widehat{({\partial_{\theta}G}_{i,\ell})}_{j}=e_{i}^{\prime}\partial_{\theta}G_{j}(\theta)e_{\ell}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/7}}\right).$
Notice that
$\hat{Q}_{j}$ $=Q_{j}+O\left(\max\{\|\hat{G}-G\|,\|\widehat{({\partial_{\theta}G}_{i,\ell})}_{j}-{({\partial_{\theta}G}_{i,\ell})}_{j}\|\}\right)=Q_{j}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/7}}\right).$
By an argument similar to the above,
$\hat{\kappa}_{\tau}$ $=\left(\tau-\frac{1}{2}\right)\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{n}\}}{2h_{n}}ZW^{\prime}G^{-1}Z\right]$
$+\sum_{j=1}^{k}\left(\tau-\frac{1}{2}\right)\mathbb{E}_{n}\left[\frac{1\{Y\leq W^{\prime}\hat{\theta}_{\tau}+h_{n}\}-1\{Y\leq W^{\prime}\hat{\theta}_{\tau}-h_{n}\}}{2h_{n}}(e^{\prime}_{j}Z_{j})ZW^{\prime}\right](\hat{G}^{-1}-G^{-1})e_{j}$
$=\kappa_{\tau}+O_{p}\left(\frac{1}{\sqrt{n}}\right)+O_{p}\left(\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right)+o_{p}\left(\frac{\log h_{n}}{n^{5/6}h_{n}}\right)+O(h_{n}^{2})+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right).$
where the last term is based on the fact
$\|\hat{G}^{-1}-G^{-1}\|\leq\|\hat{G}^{-1}\|_{2}\|G^{-1}\|_{2}\|\hat{G}-G\|.$ (48)
So the overall rate is
$\max\left\{n^{-1/2},\,\frac{\sqrt{\log n}}{n^{2/5}},\,\frac{\log h_{n}}{n^{5/6}h_{n}},\,h_{n}^{2},\,\frac{\sqrt{\log h_{n}}}{\sqrt{nh_{n}}}\right\}.$
The optimal bandwidth is $h_{3,n}\propto n^{-1/5}$ with
$\hat{\kappa}_{\tau}=\kappa_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{2/5}}\right).$
By the CLT and the equicontinuity of the relevant sample moment functions implied by Lemma A.3,
$\hat{\Omega}_{\tau}$ $\triangleq\widehat{Var}[Z(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)]$
$=\mathbb{E}_{n}[(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)ZZ^{\prime}]$
$-\mathbb{E}_{n}[Z(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)]\mathbb{E}_{n}[Z^{\prime}(1\{Y\leq W^{\prime}\hat{\theta}_{\tau}\}-\tau)]$
$=\Omega_{\tau}+O_{p}\left(\frac{1}{\sqrt{n}}\right).$
Appendix C Illustration of approximate bias formula in univariate case
Suppose we are interested in estimating the $\tau$-quantile of a uniformly distributed outcome variable $Y$. This is a special case of the general framework with $W=Z=1$, $f_{Y}(y)=1\{0\leq y\leq 1\}$.
Note that, under the maintained assumptions, the true parameter $\theta_{\tau}$ has an equivalent alternative definition as a solution to
$\mathbb{E}[(1\{-Y\leq W^{\prime}(-\theta_{\tau})\}-(1-\tau))Z]=0.$
As a result, there are two ways of defining an estimator: as a minimizer of $|\hat{g}_{\tau}(\theta)|$ or as a minimizer of $|\hat{g}_{\tau}^{*}(-\theta)|$, where
$\hat{g}_{\tau}(\theta)=\mathbb{E}_{n}(1\{Y\leq\theta\}-\tau),$
$\hat{g}_{\tau}^{*}(-\theta)=\mathbb{E}_{n}1\{-Y\leq-\theta\}-(1-\tau).$
The derivatives of the population moment conditions $g_{\tau}(\theta)=g_{\tau}^{*}(-\theta)=0$ are $G=1$, $\partial_{\theta}G=0$ and $G^{*}\triangleq\partial_{\theta}g_{\tau}^{*}(-\theta)=-1$, $\partial_{\theta}G^{*}=0$, respectively. In either case, the closure of the argmin set will be $[Y_{(k)},Y_{(k+1)}]$, where $k\triangleq\lfloor\tau n\rfloor$. If the fractional part $\{\tau n\}\triangleq\tau n-\lfloor\tau n\rfloor\leq\frac{1}{2}$, a minimizer of $|\hat{g}_{\tau}(\theta)|$ ($|\hat{g}_{\tau}^{*}(-\theta)|$) is the order statistic $Y_{(k)}$ ($Y_{(k+1)}$, respectively); if $\{\tau n\}\geq\frac{1}{2}$, a minimizer of $|\hat{g}_{\tau}(\theta)|$ ($|\hat{g}_{\tau}^{*}(-\theta)|$) is $Y_{(k+1)}$ ($Y_{(k)}$, respectively). Of course, on the real line $\mathbb{R}^{1}$, all norms $\|\cdot\|_{p},\leavevmode\nobreak\ p\in[1,\infty]$, coincide with the absolute value ||$\cdot$|| ⋅ |.
In this simple example, formula (8) yields asymptotic bias expansions
$\mathbb{E}Y_{(k)}-\tau=\frac{k-\tau n}{n}+\frac{1}{n}\left(\frac{1}{2}-\tau\right)-\frac{1}{2n}+o\left(\frac{1}{n}\right)=-\frac{\{\tau n\}}{n}-\frac{\tau}{n}+o\left(\frac{1}{n}\right),$ (49)
$\mathbb{E}Y_{(k+1)}-\tau=\frac{k-\tau n}{n}+\frac{1}{n}\left(\frac{1}{2}-\tau\right)+\frac{1}{2n}+o\left(\frac{1}{n}\right)=-\frac{\{\tau n\}}{n}+\frac{1-\tau}{n}+o\left(\frac{1}{n}\right).$ (50)
The exact bias formulas are given by (e.g., Ahsanullah et al., 2013)
$\mathbb{E}Y_{(k)}-\tau=\frac{k}{n+1}-\tau=-\frac{\{\tau n\}}{n+1}-\frac{\tau}{n+1},$
$\mathbb{E}Y_{(k+1)}-\tau=\frac{k+1}{n+1}-\tau=-\frac{\{\tau n\}}{n+1}+\frac{1-\tau}{n+1}.$
Comparing these formulas with the asymptotic formulas (49) and (50), we see that they indeed coincide up to $O\left(n^{-2}\right)$. Figure 1 in the main text illustrates the exact and the second-order bias formula (scaled by $n$) for $n=10$.
Appendix D Exact QR and IVQR algorithms
First consider a linear programming (LP) implementation of the QR regression (2) (Koenker, 2005, Section 6.2):
$\min_{\theta,r,s}\tau\bm{\iota}^{\prime}r+(1-\tau)\bm{\iota}^{\prime}s$
$s.t.\leavevmode\nobreak\ \leavevmode\nobreak$ $\varepsilon_{i}=r_{i}-s_{i}=Y_{i}-W_{i}^{\prime}\theta,\,\,i=1,\dots,n,$
$r_{i}\geq 0,s_{i}\geq 0,\,\,i=1,\dots,n.$
Here $\bm{\iota}$ is an $(n\times 1)$ vector of ones. This formulation allows us to apply LP solvers like Gurobi to obtain the exact minimum in (2).
Next consider the exact estimator for the IVQR case,
$\hat{\theta}_{\tau,1}=\operatorname{argmin}_{\theta\in\Theta}||\hat{g}_{\tau}(\theta)||_{1}.$
The underlying optimization problem can be equivalently reformulated as a mixed integer linear program (MILP) with special ordered set (SOS) constraints,
$\min_{e,\theta,r,s,t}\bm{\iota}^{\prime}t$
$s.t.\leavevmode\nobreak\ \leavevmode\nobreak$ $\varepsilon_{i}=r_{i}-s_{i}=Y_{i}-W_{i}^{\prime}\theta,\,\,i=1,\dots,n,$
$(r_{i},e_{i})\in SOS_{1},\,\,i=1,\dots,n,$
$(s_{i},1-e_{i})\in SOS_{1},\,\,i=1,\dots,n,$
$r_{i}\geq 0,s_{i}\geq 0,\,\,i=1,\dots,n,$
$e_{i}\in\{0,1\},\,\,i=1,\dots,n,$
$-t_{l}\leq{Z}_{l}^{\prime}(e-\tau\bm{\iota})\leq t_{l},\,\,l=1,\dots,d.$
where ${Z}_{l}$ is an $n\times 1$ vector of realizations of instrument $l$. All constraints except the last one coincide with the ones derived by Chen and Lee (2018) in Appendix C.1 (we also omit the redundant constraint $r_{i}+s_{i}>0$, which is implied by the two $SOS_{1}$ constraints). The last constraint ensures that the objective function is the $\ell_{1}$ norm of the just identifying moment conditions.
Remark 3.
We also considered the “big-M” formulation while performing the Monte Carlo analyses. The big-M formulation has certain computational advantages, although the arbitrary choice of tuning parameters may result in sub-optimal solutions. This problem is more prominent for tail quantiles. Since the big-M formulation does not guarantee exact solutions, consistent with our theory, the choice of tuning parameters may affect the asymptotic bias. We prefer the above SOS formulation because it does not depend on tuning parameters as the big-M MILP/MIQP formulations in Chen and Lee (2018) and Zhu (2019). [^9] These papers pick the value of the tuning parameter $M$ as a solution to a linear program that in turn depends on the choice of an arbitrary box around a linear IV estimate. This is problematic if there is a lot of heterogeneity in the coefficients across quantiles. Moreover, in linear models with heavy tailed residuals, the linear IV estimator is not consistent.
Appendix E Stochastic expansion of 1-step corrected IVQR estimators
In the main text, we focus in classical QR and exact IVQR estimators. As shown in the following corollary, the results in Theorem 1 can be used to obtain a uniform BK expansion for general IVQR estimators after a feasible 1-step correction.
Corollary 1.
Suppose that Assumptions 13 hold. Consider any estimator $\hat{\theta}_{\tau}$ such that $\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\theta_{\tau}\|=O_{p}\left(n^{-1/2}\right)$. Then
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{\theta}_{\tau}-\hat{G}^{-1}\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{\xi}_{\tau}\|=O_{p}\left(\frac{\sqrt{\log n}}{n^{3/4}}\right),$
where $\hat{G}$ is defined in (9).
Proof.
By Lemma A.2 applied to $\hat{\theta}_{\tau}$, uniformly in $\tau\in[\varepsilon,1-\varepsilon]$,
$\hat{\theta}_{\tau}-G^{-1}\hat{g}_{\tau}(\hat{\theta}_{\tau})=\hat{\xi}_{\tau}+O_{p}\left(\frac{\sqrt{\log n}}{n^{3/4}}\right).$
Under the maintained assumptions, Lemma 1 implies
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{G}(\hat{\theta}_{\tau})-G(\theta_{\tau})\|=O_{p}\left(n^{-2/5}\sqrt{\log n}\right).$
By Lemma A.1, $\partial G(\theta)$ is bounded uniformly over $\theta\in\Theta$. Then, by Assumption 1.2 and continuity of the minimal eigenvalue function, the eigenvalues of $G(\theta)$ are bounded away from zero on $\theta\in\Theta$. Therefore, the derivative of the inverse matrix function, $F(A)\triangleq A^{-1}$, is uniformly bounded over $G(\theta_{\tau})$ for $\tau\in[\varepsilon,1-\varepsilon]$. Hence, by the element-wise Taylor expansion of $F$ at $G(\theta_{\tau})$,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{G}^{-1}(\hat{\theta}_{\tau})-G^{-1}(\theta_{\tau})\|=O\left(\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{G}(\hat{\theta}_{\tau})-G(\theta_{\tau})\|\right)=O_{p}\left(n^{-2/5}\sqrt{\log n}\right).$
By Lemma A.3,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau})-\hat{g}_{\tau}(\theta_{\tau})\|=O_{p}\left(\frac{\sqrt{\log n}}{n^{3/4}}\right).$ (51)
By Donsker’s theorem,
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\theta_{\tau})\|=O_{p}\left(\frac{1}{n^{1/2}}\right),$
so, by the triangular inequality and (51),
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{g}_{\tau}(\hat{\theta}_{\tau})\|=O_{p}\left(\frac{1}{n^{1/2}}\right).$
Then
$\sup_{\tau\in[\varepsilon,1-\varepsilon]}\|\hat{G}^{-1}(\hat{\theta}_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau})-G^{-1}(\theta_{\tau})\hat{g}_{\tau}(\hat{\theta}_{\tau})\|=O_{p}\left(n^{-1/2}n^{-2/5}\sqrt{\log n}\right),$
which concludes the proof. ∎
Appendix F Additional figures
In this section, we present additional simulation results. To explore the impact of the strength of the instruments, we consider two additional DGPs.
[…]
In DGP7, the instrument is weaker than in DGP1, and in DGP8, the instrument is stronger than in DGP4.
Figure 7: Bias (multiplied by $n$) before and after correction, IV strength grows from left to right
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP7, DGP4, and DGP8 (ordered in terms of instrument strength). All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
Figure 8: Bias (multiplied by $n$) before and after correction for DGP1, different sample sizes
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for classical QR without bias correction (blue dots), QR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and QR with infeasible bias correction (gold dashed line) for DGP1 with $n\in\{50,100,200\}$. All results are based on 5,000 simulation repetitions.
Figure 9: Bias (multiplied by $n$) before and after correction for DGP4, different sample sizes
Notes: The panels display the bias (multiplied by $n$) of the intercept and the slope for IVQR (implemented via the MILP formulation in Appendix D) without bias correction (blue dots), IVQR with feasible bias correction based on the rule-of-thumb bandwidth (gold squares), and IVQR with infeasible bias correction (gold dashed line) for DGP4 with $n\in\{50,100,200\}$. All results are based on 5,000 simulation repetitions. The infeasible bias correction is based on the feasible formula applied to a simulated sample of 10,000,000 observations.
Figure 10: RMSE comparison of raw and bias-corrected estimators
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP2, classical QR and (b) DGP5, IVQR. All results are based on 5,000 simulation repetitions.
Figure 11: RMSE comparison of raw and bias-corrected estimators
Notes: The panels compare the RMSE for estimators without bias correction (blue), with infeasible bias correction (grey) and with feasible bias correction based on the rule-of-thumb bandwidth choice (gold) for (a) DGP3, classical QR and (b) DGP6, IVQR. All results are based on 5,000 simulation repetitions.
Figure 12: Mean absolute deviations (MAD) of raw and bias-corrected estimators
Notes: The panels compare the Mean Absolute Deviation for the estimator without bias correction (blue) and with feasible bias correction (gold) for (a) DGP1, classical QR and (b) DGP4, IVQR. All results are based on 5,000 simulation repetitions.
Figure 13: Confidence interval coverage before and after correction for $n=100$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP2 (classical QR) and DGP5 (IVQR). All results are based on 5,000 simulation repetitions.
Figure 14: Confidence interval coverage before and after correction for $n=100$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP3 (classical QR) and DGP6 (IVQR). All results are based on 5,000 simulation repetitions.
Figure 15: Confidence interval coverage before and after correction for $n=200$
Notes: The panels display the coverage probability of the $90\%$ confidence intervals for the intercept and the slope without bias correction (blue dots) and with the feasible bias correction based on the rule-of-thumb bandwidth choice (gold squares) for DGP1 (classical QR) and DGP4 (IVQR). All results are based on 5,000 simulation repetitions.